### v1

# Evaluating Human-AI Collaboration: A Review and Methodological Framework

Abstract
The use of artificial intelligence (AI) in working environments with individuals, known as Human-AI Collaboration (HAIC), has become essential in a variety of domains, boosting decision-making, efficiency, and innovation. Despite HAIC’s wide potential, evaluating its effectiveness remains challenging due to the complex interaction of components involved. This paper provides a detailed analysis of existing HAIC evaluation approaches and develops a fresh paradigm for more effectively evaluating these systems. Our framework includes a structured decision tree which assists to select relevant metrics based on distinct HAIC modes (AI-Centric, Human-Centric, and Symbiotic). By including both quantitative and qualitative metrics, the framework seeks to represent HAIC’s dynamic and reciprocal nature, enabling the assessment of its impact and success. This framework’s practicality can be examined by its application in an array of domains, including manufacturing, healthcare, finance, and education, each of which has unique challenges and requirements. Our hope is that this study will facilitate further research on the systematic evaluation of HAIC in real-world applications.
Keywords Human-AI Collaboration (HAIC)  $\cdot$ HAIC Evaluation  $\cdot$ Symbiotic AI Systems  $\cdot$ AI Metrics  $\cdot$ AI Evaluation Framework
1 Introduction
The fast advancement of Artificial Intelligence (AI) has fundamentally transformed several domains, blurring the distinctions between human and machine roles and giving rise to new interactions [1]. The notion of Human-AI Collaboration (HAIC) has emerged as a pivotal framework, as evidenced by the research conducted by [2] and [3]. These studies identify that HAIC has significant potential for enhancing decision-making, increasing efficiency, and fostering innovation. They therefore aim at fostering and facilitating effective cooperation between people and AI systems to achieve common goals. On the other hand, assessing the benefits of this cooperation presents its own challenges, owing to the multiple complex factors involved.
Traditional Human-Machine evaluation frameworks have primarily focused on quantifiable metrics such as task performance, response time, and user satisfaction within human-machine systems [4, 5]. These frameworks typically view the machine as a tool and assess how effectively humans can control or interact with this tool to achieve desired outcomes.
Human-AI Collaboration (HAIC) represents a significant change in which humans and AI work together in a more integrated and mutually beneficial manner. HAIC focuses on more than just human dominance over machines; it emphasizes a collaborative partnership where individuals and AI systems actively contribute to shared results. The significance of this partnership has been highlighted in several domains, such as manufacturing [6, 7], financial services [8, 9], healthcare [10, 11], education [12], and others, demonstrating notable advancements in both technological innovation and practical implementation [13, 14].
HAIC is defined by the process of making decisions together, learning from each other, and adapting, which requires a more comprehensive and nuanced evaluation approach [15, 16]. To effectively evaluate HAIC, it is crucial to consider not only the traditional metrics of task success but also the quality of interactions within the team, the processes by which decisions are made, and crucial ethical aspects such as the transparency of algorithms and the mitigation of biases [17]. This involves assessing how AI systems influence human decisions, how humans and AI adapt to each other’s capabilities, and how ethical considerations are integrated into collaborative processes. In this paper, we aim at addressing these evolving needs by critically analyzing current evaluation methods in human-AI interaction and proposing a new framework designed to capture the dynamic and reciprocal nature of human-AI collaborations. This framework seeks to recognize and measure both the individual and joint contributions of humans and AI to collaborative outcomes, thus providing a more comprehensive understanding of HAIC effectiveness.
The contributions of this paper are therefore the following.
  • Overview of the Literature: Extending the foundational work of Madni [2] and Dubey et al. [3], we provide a review of HAIC evaluation methods, identifying gaps and research opportunities in the existing literature.
  • Novel Evaluation Framework: Inspired by the pioneering frameworks of Rouse et al.[18] and Parasuraman et al.[19, 4], we propose a new, adaptable evaluation framework for HAIC that incorporates both quantitative and qualitative measures, emphasizing ethical considerations and providing a flexible structure for diverse HAIC applications.
  • Domain-Specific Insights: Through detailed examination of applications in healthcare, finance, and education, as informed by the studies of Tschandl et al.[10], Buckley and Lui[8, 9], and Holstein et al. [12], we demonstrate the applicability and customization of our framework to meet domain-specific needs.
Structure of the Paper
Following this introduction, Section 2 discusses the fundamentals of Human-AI Collaboration, including various modes of interaction. Section 3 reviews the existing literature on HAIC, highlighting advancements in evaluation frameworks and tools, as well as the diversity and challenges in domain-specific evaluations. Section 4 presents a structured framework for evaluating HAIC, detailing factors, subfactors, and metrics for assessment. Section 5 applies this evaluation framework to real-world domains such as manufacturing, healthcare, finance, and education. Section 6 addresses the unique challenges of evaluating creative and linguistic AI, including large language models and generative AI in the arts. Section 7 discusses the limitations in investigating behavioral factors. Finally, Section 8 discusses the implications of the HAIC evaluation framework for future research and practice.
2 Understanding the Fundamentals of Human-AI Collaboration
2.1 From Human-Machine Interaction to Human-AI Collaboration
Given the rapid pace of advancements in Human-AI Collaboration (HAIC), a comprehensive understanding of the current landscape is essential. While foundational research in Human-Machine Interaction (HMI) provides valuable insights into human-technology interactions, it often falls short when applied to HAIC evaluations. Traditional HMI methods, as exemplified in the works of Nielsen [20], Shneiderman [21], and Norman [22], primarily focus on assessing the usability and effectiveness of tools or interfaces designed to augment human performance. These methods often rely on metrics such as task completion time, error rates, and user satisfaction to gauge the success of human-machine interactions. However, these metrics are not always sufficient to capture the complexities of HAIC, where the relationship between humans and AI is more dynamic and collaborative. For example, traditional metrics may not adequately assess the level of trust a human places in an AI system’s recommendations or the effectiveness of shared decision-making processes where both the human and AI contribute to the final outcome [23]. Additionally, these metrics may not fully capture the adaptability of AI systems, which is crucial for HAIC as the AI needs to learn and adjust its behavior based on human input and changing circumstances [24]. To effectively evaluate HAIC, we must move beyond the limitations of traditional HMI approaches. We need to consider not just the performance of the individual components (human and AI) but also the quality of their collaboration, the fluidity of their interactions, and the achievement of shared goals. This necessitates a broader range of metrics and methodologies that can capture the nuances of these complex partnerships.
2.2 Elements of HAIC
Human-AI Collaboration (HAIC) is the cooperative partnership between individuals and artificial intelligence (AI) systems in order to accomplish common goals.
Several key elements define the landscape of HAIC:
  • Tasks: HAIC systems can be designed to tackle a wide array of tasks, from complex decision-making to creative endeavors and knowledge work. The nature of these tasks often dictates the level and type of collaboration required between humans and AI [25, 26]. For example, in the medical field, an AI system can analyze medical images to identify potential anomalies, while a radiologist interprets the results and makes a final diagnosis.
  • Goals: HAIC is driven by shared goals, which can be both individual and collective. Individual goals might encompass AI objectives like improving efficiency and accuracy or human aims like skill enhancement and knowledge acquisition. Collective goals focus on achieving overarching objectives that benefit from the combined capabilities of both humans and AI [27]. For example, in a healthcare setting, an AI system could be designed to examine medical images and identify initial indications of illnesses like cancer (individual goal). The human radiologist review’s the the AI’s findings considers patient history, and makes the final diagnosis (individual goal). The collective goal in this scenario is to improve diagnostic accuracy and ensure timely and effective patient care, thereby enhancing overall health outcomes.
  • Interaction: The success of HAIC heavily relies on effective communication and feedback mechanisms between humans and AI. The quality of this interaction determines how well each party understands the other’s intentions, capabilities, and limitations [17, 28]. For instance, an AI chatbot for customer service needs to understand and respond to customer queries effectively, adapting its communication style based on user feedback.
  • Task Allocation: Assigning tasks appropriately based on the strengths and weaknesses of both humans and AI is critical. Dynamic task allocation, where responsibilities shift based on real-time needs, is often a hallmark of successful HAIC [3]. A good example is in autonomous vehicles, where AI systems may be responsible for navigation under normal conditions, but the human driver takes control when faced with unexpected situations.
Figure 1 visually represents the key elements of HAIC, which are important for developing effective evaluation frameworks. By clearly defining the nature of the collaboration, the intended goals, and the dynamics of interaction, we can then create metrics and assessments that accurately measure the success and impact of HAIC systems.
Figure 1: Core elements of Human-AI Collaboration (HAIC) and examples of their application.
2.3 Modes of Human-AI Collaboration
One can use any of the HAIC elements, i.e., tasks, goals, types of interaction, and task allocation to differentiate between different HAIC methods. For the purposes of evaluation, however, perhaps the most useful distinction is in terms of allocation [29]. We identify three modes, namely AI-centric, Human-centric, and Symbiotic, which are further described below.
Human-Centric Mode
In the Human-centric mode, humans retain the primary decision-making authority, utilizing AI as an augmentative tool to enhance human capabilities without superseding the human role. This mode values human intuition and oversight, with AI employed to manage data-intensive or repetitive tasks. It optimizes the unique cognitive abilities of humans for strategic and ethical decision-making while leveraging AI to augment human capabilities.
Supporting this mode, Sutthithatip [30] highlights the use of Explainable AI (XAI) in high-stakes sectors like aerospace to demystify complex AI operations, thereby enhancing human decision-making. Additionally, How et al. [31] describe the integration of AI in sustainable development, which facilitates the involvement of non-experts in data-driven policy making, illustrating the supportive role of AI in enhancing human capabilities.
Symbiotic Mode
The Symbiotic mode represents a balanced partnership where humans and AI systems collaborate closely, mutually enhancing each other’s capabilities. This mode is characterized by a two-way interaction, shared decision-making, and a continuous exchange of feedback, aiming to achieve collective goals through a synergistic relationship. This mode is particularly relevant in complex tasks where both human intuition and AI’s computational capabilities are essential for optimal outcomes.
Studies like those by Cesta et al. [32] explore the symbiotic integration of AI in operational settings, enhancing the collaborative dynamics between humans and robots. Moreover, Jokinen & Watanabe [33] emphasize the crucial role of trust in these interactions, which is fundamental for a successful symbiotic relationship. Mahmud et al. [34] and Sowa et al. [35] further illustrate how symbiotic HAIC can revolutionize creative fields and managerial roles by combining human creativity with AI’s analytical strengths to foster innovation and improve decision-making.
A key example of the symbiotic mode is the Learning to Defer paradigm. This paradigm designs AI systems to understand their own limitations, knowledge, and expertise. Whether confronted with ambiguity or circumstances in which their confidence in a decision is low, the AI system proactively defers to human judgment  [36]. Several studies have explored this concept in depth. Mozannar and Sontag (2020) [37] proposed a Bayesian framework that emphasizes the integration of human and machine predictions. It is crucial to consider potential biases in both agents. This approach aligns with a collaborative decision-making approach because it strives for optimization. It is possible to make decisions by leveraging the strengths of both humans and AI. In a similar way, Mao et al. (2023) [38] extended this concept to scenarios with multiple experts, highlighting the flexibility required in real-world HAIC systems, where multiple sources of expertise may be available. Their work underscores the adaptability of the HAIC systems’ participants, showcasing the ability to dynamically modify task allocation according to the availability and expertise of various agents.
AI-centric Mode
The AI-centric mode designates AI as the primary agent in the collaboration, where the AI system leads decision-making processes and operates with minimal human intervention. This mode often features automated interactions where the AI executes tasks independently, aimed at enhancing system capabilities and overall efficiency. Such a setup focuses on improving AI functionalities, maximizing system efficiencies, and generally involves a one-directional interaction from AI to human, primarily concentrating on tasks suited for AI’s computational abilities.
Recent research underscores the evolving role of AI in traditional human-machine settings. Liu and Zhao [39] discuss the allocation of roles in Human-Machine Collaboration (HMC), advocating for optimal strategies that seamlessly integrate intelligent systems into human-centric operations. Similarly, Wang et al. [40] examine the shift from human-human to human-AI collaboration, stressing the importance of mutual goal setting and collaborative task management to foster effective partnerships.
In this section, we looked at the fundamentals of Human-AI Collaboration, such as its main components and the various modes of operation. Evaluating HAIC is critical to ensuring that these systems are successful, efficient, and ethical.
The following part looks into the literature on HAIC evaluation, reviewing existing approaches and identifying gaps that suggest areas for further research.
3 Evaluation of Human-AI Collaboration in the Research Literature
The evaluation of Human-AI Collaboration (HAIC) is essential to understanding the efficacy of these systems, identifying areas for improvement, and ultimately unlocking their full potential. This section examines current evaluation approaches in HAIC research. While Table 1 provides an overview of a diverse sample of studies across different domains, the analysis in this section will focus on selected key references to illustrate specific evaluation methodologies and their implications.
Through this analysis of a wide array of studies, our objectives are as follows:
  • Examine the advantages and drawbacks of current evaluation methodologies.
  • Identify the distinct difficulties presented by various HAIC applications and domains.
  • To effectively address the intricate and diverse aspects of HAIC in different sectors.
{adjustwidth}
-1cm-1cm Reference Domain AI Technology/Method Qn Ql Mix Additional Info Vollmuth et al. (2023) [41] Healthcare (Oncology) Image Analysis Accuracy is measured by sensitivity, precision, recall, F1-score. Focuses on human specialists to evaluate AI and HAIC performance. Van Leeuwen et al. (2022) [42] Healthcare (Radiology) Image Analysis Emphasizes efficiency using sensitivity, specificity, and outcome metrics. Timmons et al. (2023) [43] Healthcare (Mental Health) Not specified Investigates bias through interviews and case studies. Farić et al. (2024) [44] Healthcare (Radiology) Diagnostic Support Systems Assesses early AI use via qualitative feedback from radiologists. Calisto et al. (2022) [45] Healthcare (Breast Cancer) Medical Image Analysis Focuses on human-AI interaction through qualitative methods. Rezwana et al. (2023) [46] Design Generative AI Develops interaction design framework using user feedback. Tschandl et al. (2020) [10] Healthcare (Dermatology) Image Analysis Uses quantitative success metrics in high-stakes environments. Siu et al. (2021) [47] Game-Based Reinforcement Learning Uses game performance metrics in Hanabi game as a testbed. Reddy et al. (2021) [48] Healthcare Not specified Proposes a qualitative framework focusing on ethics and user perspectives. Sharma et al. (2023) [49] Mental Health Reinforcement Learning Measures conversational empathy improvements through qualitative feedback. Yik et al. (2024) [50] General (Benchmarking) Neuromorphic Computing Provides quantitative framework for unique AI context. Rastogi et al. (2023) [51] General (Auditing) LLM Examines auditing collaboration using qualitative assessments. Zhao et al. (2023) [52] Multimodal Interaction LLM with Visual Input Studies interaction through multimodal input performance metrics. Makarius et al. (2020) [1] General (Business) Not specified Emphasizes social factors using qualitative methods. Aleid et al. (2023) [53] Healthcare (Radiology) Image Analysis Focuses on accuracy using quantitative measures. Al-Fatlawi et al. (2023) [54] Finance (Fraud) Genetic Algorithms Evaluates detection rates (true/false positives). Yang et al. (2022) [55] General (Teaming) Behavior Modeling Uses quantitative and qualitative data to model human behavior for better interaction. Aher et al. (2023) [56] Simulation-Based Large Language Models Simulates human behavior using quantitative metrics. Verheijden & Funk [57] (2023) Design Generative AI Enhances design processes through qualitative co-creation feedback. Hauptman et al. (2023) [58] General (Teaming) Adaptive AI Investigates adaptability perceptions through qualitative assessments. Sankaran et al. (2022) [59] Manufacturing Recurrent Neural Networks Measures process performance using quantitative metrics. Massaro (2022) [60] Manufacturing (Safety) Artificial Neural Networks Evaluates safety and risk through quantitative measures. Oravec (2023) [61] Education Large Language Models Discusses academic integrity using qualitative methods. El-Zanfaly et al. (2022) [62] Design/Art Physical Interface Focuses on creativity and interaction through qualitative feedback. Arias-Rosales (2022) [63] Design/Art Shape Generation Combines quantitative shape metrics with qualitative value assessments. Dziorny et al. (2022) [64] Healthcare (ICU) Decision Support Systems Examines high-stakes challenges using qualitative feedback. Vössing et al. (2022) [65] General (Forecasting) LightGBM, Neural Networks Investigates AI transparency impact using mixed trust/performance metrics. Chakravorti et al. (2022) [66] Finance Prediction Models Combines detection rates with qualitative trust assessments for AI markets. Huang & Rust (2022) [67] Marketing Machine Learning Develops collaborative AI framework using mixed methods. Sachan et al. (2024) [68] Finance Evidential Reasoning Reduces decision noise through quantitative measures. Basu et al. (2021) [69] Finance (ICO) Linear Regression Studies decision-making using mixed quantitative and qualitative data. Dikmen & Burns (2022) [70] Finance (Peer-to-Peer Lending) Explainable AI Examines trust and performance using mixed domain knowledge assessments. Ma et al. (2024) [71] General (Decision-Making) Large Language Models Proposes deliberative AI using mixed decision-making metrics. Alon-Barkat & Busuioc (2023) [72] Public Sector Regression Models Explores public sector interactions through mixed methods. Nikolopoulou (2024) [73] Education Generative AI Assesses generative AI’s impact on pedagogy using mixed methods.
Table 1: Generalized Evaluation of Human-AI Collaboration Across Sectors
3.1 Advancements in Evaluation Frameworks
Human-AI Collaboration is being judged by a more traditional metrics. These include ethical concerns, the ability to scale and the ability to adapt different domains.
Dellermann et al. (2021) [25] propose a classification system for hybrid intelligence, emphasizing the importance of understanding the nature of tasks and the dynamic relationship between humans and AI. To categorize different types of HAIC systems, their framework takes into account factors such as level of autonomy, task interdependence, and degree of human control. This classification scheme serves as a foundation for developing tailored evaluation approaches that address the unique characteristics of each type of collaboration.
The incorporation of ethical factors into evaluation frameworks is a growing trend, as evidenced by the work of Huang & Rust (2022) [67] in marketing and Nasir, Khan, & Bai (2023) [74] in healthcare. These frameworks emphasize the importance of transparency (e.g., disclosing the use of AI in decision-making), fairness (e.g., mitigating bias in algorithms), accountability (e.g., establishing clear lines of responsibility for AI-related outcomes), and human-centered design (e.g., prioritizing user needs and values) in AI implementation. Lase & Nkosi (2023) [75] further advocate for a Human-centric perspective in AI development, stressing the significance of aligning AI systems with user values and ethical principles. They propose using surveys and interviews to assess user perceptions of fairness, trust, and the impact of AI on their work.
Bojic et al. (2023) [76] highlight the need for standardized and robust evaluation practices, proposing a hierarchical framework to guide the design of reliable human evaluations of AI systems. This framework outlines a step-by-step process for defining evaluation objectives, selecting appropriate metrics (e.g., accuracy, efficiency, user satisfaction), designing experimental protocols, and analyzing data to draw valid conclusions about the effectiveness of HAIC.
3.2 Advancements in Evaluation Tools and Methodologies: Enabling Real-time Insights and Adaptive Collaboration
Technological advancements are transforming the landscape of HAIC evaluation, enabling real-time monitoring, adaptive collaboration, and deeper insights into human-AI interactions.
AI-powered analytics and real-time monitoring tools, discussed by Wang et al. (2019) [77], are enhancing the precision and responsiveness of HAIC evaluations. These tools can automatically track key performance indicators (KPIs) such as completion time, error rates, and user interactions, providing immediate feedback on collaboration efficacy and enabling timely interventions for improvement.
Advancements in interactive machine learning (iML) interfaces, explored by Saha et al. (2023) [78], are fostering grater transparency and user understanding. By providing explanations for AI-generated recommendations and allowing users to adjust system parameters, iML interfaces can improve user comprehension, trust, and satisfaction, ultimately contributing to a more effective and collaborative partnership between humans and AI. Saha et al. [78] evaluate the effectiveness of iML interfaces using a combination of user surveys, interviews, and task-based performance metrics.
3.3 Analysis of Existing Approaches
In Table 1, we show a summary of Human-AI Collaboration (HAIC) evaluation practices we examined for this analysis. While quantitative methods remain dominant, there is a growing recognition of the value of qualitative and mixed-methods approaches for capturing the complex interplay between objective performance and subjective experiences.
Quantitative Focus: Objectivity and Measurable Outcomes
Quantitative evaluations, prioritizing objective measures to gauge system performance and efficacy,including the combination of humans and AI, are prevalent in HAIC research. In the medical field, Vollmuth et al. (2023) [41] and Van Leeuwen et al. (2022) [42] use tests like sensitivity, specificity, precision, recall, and F1-score to carefully check how well AI-assisted tools in oncology and radiology can make diagnoses. To compare how well AI can find fraud in the financial world, Al-Fatlawi et al. (2023) [54] use detection rates (true positives and false positives), and Sachan et al. (2024) [68] use quantitative measures to reduce decision noise in financial underwriting. Manufacturing studies, such as those by Sankaran et al. (2022) [59] and Massaro (2022) [60], also rely heavily on quantitative metrics to assess process performance and safety compliance, respectively. Researchers have developed quantitative frameworks to assess performance in diverse fields such as game-based AI (Siu et al., 2021) [47] and neuromorphic computing (Yik et al., 2024) [50] based on specific domain-relevant metrics.
While these quantitative approaches provide valuable insights into system performance and measurable outcomes, their focus on objective data often overlooks the nuanced human factors that influence the adoption and success of HAIC systems.
Qualitative Insights: Illuminating the Subjective Dimension
To complement the quantitative focus, a growing number of studies are embracing qualitative methods to delve into the subjective experiences of users and stakeholders. Timmons et al. (2023) [43] investigate biases in mental health AI through interviews and case studies, shedding light on potential ethical concerns and user perceptions. Their study reveals potential ethical concerns, such as the perpetuation of biases and the varying levels of trust users place in AI systems. Rezwana et al. (2023) [46] and Verheijden & Funk (2023) [46] use qualitative assessments in the design field to understand the impact of generative AI on creativity and design co-creation processes. Through focus group discussions and in-depth interviews with designers, these studies highlight how generative AI tools influence creative workflows. They discovered that while AI can enhance creative output by providing novel ideas, it also raises concerns about authorship and originality, which are critical to the design profession.
Sharma et al. (2023) [49] look at how conversational empathy improves in mental health applications using qualitative measures. Reddy et al. (2021) [48] suggest a qualitative framework for evaluating the use of AI in healthcare, with a focus on ethical issues and user perspectives. Their framework includes structured interviews with healthcare providers and patients, and thematic analysis to identify key concerns such as data privacy, trust in AI diagnoses, and the perceived impact on the patient-provider relationship. Their study underscores the necessity for AI systems to align with ethical standards and user expectations to be effectively integrated into healthcare practices.
By capturing the nuanced perspectives of users and stakeholders, these qualitative studies offer a rich understanding of the human factors that shape the adoption and impact of HAIC. However, the subjective nature of qualitative data can pose challenges in terms of generalizability and require substantial resources for data collection and analysis.
The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
The limitations of purely quantitative or qualitative approaches have spurred the adoption of mixed-methods evaluations, aiming to capture both objective performance and subjective experiences in a comprehensive manner.
Domain-Specific Applications
In the realm of design and art, Arias-Rosales (2022) [63] combines quantitative shape generation metrics with qualitative assessments of perceived value by designers. Their study uses quantitative metrics to measure the efficiency and novelty of AI-generated shapes, while also conducting interviews and surveys with designers to gather insights on the aesthetic value and usability of these shapes. This approach revealed that while AI-generated designs were technically innovative, the perceived value varied significantly among designers, highlighting the importance of subjective aesthetic judgment in evaluating creative outputs.
Similarly, in finance, Chakravorti et al. (2022) [66] integrate quantitative detection rates with qualitative assessments of trust and perception to evaluate the efficacy of AI prediction markets. They employ quantitative methods to analyze the accuracy and reliability of AI predictions in financial markets, alongside qualitative interviews and surveys with traders and financial analysts. The study found that despite high detection rates, the trust in AI predictions varied, with qualitative data uncovering concerns about the transparency and interpretability of AI algorithms.
Generalizable Frameworks
Yang et al. (2022) [55] look into how AI can model human behavior using both quantitative and qualitative data. They use quantitative metrics such as task completion rates and error frequencies to evaluate AI performance, while also gathering qualitative data through user interviews and behavioral observations. Vössing et al. (2022) [79] study includes quantitative measures of system accuracy and user performance metrics, coupled with qualitative trust assessments obtained from user interviews and surveys. The combination of these methods revealed that transparency significantly impacts user trust and acceptance, even when the AI system’s performance is objectively high.
Huang & Rust (2022) [67] propose a mixed-methods framework for collaborative AI in marketing, acknowledging the importance of both performance and ethical considerations.These mixed-methods techniques provide the possibility for a comprehensive and contextual evaluation of HAIC, recognizing the intricate relationship between objective measurements and subjective perceptions. They employ quantitative analytics to measure marketing campaign performance metrics, such as click-through rates and conversion rates, while also using qualitative methods like focus groups and in-depth interviews to assess ethical concerns and user satisfaction. This dual approach provides a comprehensive evaluation of the AI’s effectiveness and its alignment with ethical standards.
3.4 The domain-specific challenges necessitate a tailored evaluation
The application domains for HAIC are diverse, each presenting unique challenges that necessitate tailored evaluation approaches. These challenges stem not only from the technical complexities of integrating AI into different workflows, but also from the varying priorities and expectations of stakeholders in different sectors. As a result, the evaluation methods employed across HAIC research vary significantly, reflecting the need to adapt assessment strategies to the specific context of each domain.
Healthcare: Prioritizing Accuracy, Safety, and Ethical Considerations
In the healthcare sector, ensuring diagnostic accuracy and patient safety is paramount. To find out how well AI-assisted tools work in oncology and radiology, studies like Vollmuth et al. (2023) [41] and Van Leeuwen et al. (2022) [42] use numbers like sensitivity, specificity, and outcome improvements. However, the high-stakes nature of medical decisions and the potential for unintended consequences necessitate additional layers of evaluation. Qualitative approaches like interviews and focus groups (Farić et al., 2024 [14]; Calisto et al., 2022 [45]) can capture user perceptions, identify potential biases (Timmons et al., 2023 [43]), and ensure that AI systems are implemented in a way that aligns with ethical principles and patient values (Reddy et al., 2021 [48]).
Creative Arts: Balancing Creativity, User Experience, and Ethical Impact
In the creative arts, HAIC evaluation often requires a different approach. While quantitative metrics can measure the technical performance of AI systems, such as the speed or accuracy of generating artistic outputs, qualitative methods such as interviews, user feedback, and case studies are crucial for comprehending the influence of AI on creativity, collaboration, and the overall design process (Rezwana et al., 2023 [46]; Verheijden & Funk, 2023 [57]). Additionally, ethical considerations, such as the potential for AI to perpetuate biases or undermine the role of human creativity, necessitate careful evaluation through qualitative assessments and critical analysis. Moreover, the cooperative aspect of HAIC plays a crucial role in enhancing human creativity, since AI’s data-driven insights stimulate inventive solutions [80].
Finance: Mitigating Risks, Ensuring Trust, and Balancing Quantitative and Qualitative Insights
The financial sector employs a mix of quantitative and qualitative methods to address HAIC’s unique challenges in this domain. Quantitative metrics like detection rates and false positives are crucial for evaluating the effectiveness of AI in fraud detection (Al-Fatlawi et al., 2023 [54]). However, understanding the impact of AI on decision-making processes, investor trust, and market stability requires qualitative insights through surveys, interviews, and case studies (Chakravorti et al., 2022 [66]). Mixed-methods approaches, which combine quantitative and qualitative data, offer a more comprehensive understanding of the complex interactions between humans, AI, and financial markets.
In summary, the review of evaluation approaches demonstrates significant variation in collaboration methods, domain-specific challenges, and methodological approaches. However, one significant obstacle remains: the lack of a comprehensive, mixed-methods evaluation framework that covers a wide range of industries. This gap restricts our ability to completely evaluate HAIC’s effectiveness and maximize its potential across domains.
The Path Forward: A Unified Evaluation Framework to Bridge the Gap in HAIC Assessment
The analysis of existing approaches to HAIC evaluation reveals a landscape marked by diversity and fragmentation. Quantitative, qualitative, and mixed-methods approaches all give useful information, but the lack of a single, flexible framework makes it harder to get a full picture of how well HAIC works and what it can do in different fields.
This gap is particularly evident when considering the multifaceted nature of HAIC. Quantitative metrics, while essential for measuring objective performance, often fail to capture the nuances of human-AI interaction dynamics, such as trust, satisfaction, and ethical considerations. Qualitative methods, on the other hand, can provide rich insights into these subjective experiences but may lack the rigor and generalizability of quantitative data.
Therefore, it is imperative to develop a comprehensive, mixed-methods evaluation framework. This framework should be adaptable across a broad spectrum of industries, accommodating the unique characteristics and challenges of each while providing a cohesive approach to assessing the effectiveness, challenges, and opportunities of HAIC. By integrating objective performance metrics with subjective experiences, such a framework can offer a more nuanced understanding of how HAIC systems perform in real-world contexts and how they impact various stakeholders.
This unified framework would not only address the current methodological fragmentation but also facilitate cross-study comparison, enhance the generalizability of findings, and promote standardized evaluation practices. It would enable researchers, practitioners, and policymakers to make informed decisions based on a more holistic understanding of the complex interactions between humans and AI.
In healthcare, this framework could integrate measures of diagnostic accuracy with assessments of user trust and ethical considerations, ensuring that AI-assisted tools are both effective and aligned with patient values. In creative fields, it could incorporate quantitative metrics of creative output alongside qualitative assessments of the impact on artistic processes and collaboration. In finance, it could combine quantitative measures of risk and return with qualitative assessments of trust, transparency, and the impact on decision-making.
By addressing the limitations of existing approaches and providing a standardized yet adaptable framework, we can unlock the full potential of HAIC across diverse domains. This would enable us to harness the power of AI in ways that augment human capabilities, foster innovation, and ultimately contribute to societal well-being. The development of such a framework is not merely an academic pursuit but a critical step towards ensuring the responsible and impactful integration of AI into human workflows.
4 A Structured Framework for Human-AI Collaboration
4.1 Factors of Identifying the Human-AI Collaboration
Based on the previous discussion (Section 2) the evaluation of HAIC is centered around three primary factors: Goals, Interaction, and Task Allocation. These factors are critical for understanding and enhancing HAIC systems. While tasks themselves are fundamental to HAIC, they are considered the context within which collaboration occurs, rather than a core factor of the evaluation framework. The nature of the task will influence how the other three factors are approached, but the tasks themselves are not directly evaluated by this framework.The literature highlights the importance of these factors in evaluating HAIC.
Research by Li et al. [81] and Vossing et al. [65] underscores the significance of evaluating goals through metrics such as the learning curve and prediction accuracy. Studies by Hinsen et al. [82] and Amershi et al. [83] emphasize the need for user friendly communication methods and clear interaction mechanisms for successful HAIC. Additionally, Xiong et al. [84] stress the importance of measuring query efficiency and expertise utilization for effective task allocation.
Following the identification of these primary factors, we further explore the subfactors and metrics that can be used to evaluate each factor. These subfactors and metrics are designed to provide a comprehensive understanding of the effectiveness, efficiency, and usability of HAIC systems. The structured framework presented here serves as a guide for researchers, practitioners, and policymakers to evaluate and optimize HAIC systems across various domains and applications.
4.2 Subfactors & Metrics of the Framework
To promote further analysis and evaluation of the different types of interactions in HAIC, the three primary factors are further analyzed into measurable subfactors, which can be evaluated using a set of metrics. These are presented in Table 2. Following we provide a detailed description of the subfactors and metrics associated with each primary factor. Please note that Table 3 summarizes all metrics discussed in this section, including their formal definition.
{adjustwidth}
-1cm-1cm Primary Factor Subfactor Human-AI Collaboration Metrics for Evaluation Goals Individual Goals AI’s Goal: Learn Efficiently, Increase Accuracy. Human’s Goal: Provide Effective Teaching, Achieve Task Objectives. * Learning Curve (rate of improvement over time), * Prediction Accuracy (percentage of correct predictions), * Teaching Efficiency (time/resources spent per unit of learning) [81, 85, 86]. Collective Goals Shared Goal: Maximizing the overall performance of the AI system. * Overall System Accuracy (percentage of correct outputs), * Objective Fulfillment Rate (percentage of tasks successfully completed) [65, 87]. Interaction Communication Methods Query and Response Mechanism. * Ease of Use (user satisfaction with communication methods), * Clarity of Communication (understandability of queries and responses) [82, 83] Feedback Mechanisms Human feedback on AI queries, AI adjustments based on feedback. * Frequency of Feedback (number of feedback instances), * Feedback Quality (relevance and specificity of feedback), * Feedback Impact (measured change in AI performance due to feedback) [83, 88, 82] Adaptability AI’s ability to adapt query strategy and learning based on human input. * Strategy Adaptation Effectiveness (improvement in query selection based on feedback), * Adaptability Score (composite metric of flexibility and responsiveness) [82, 89, 90] Trust and Safety Trust in AI system, ensuring user safety. * Trust Score (level of user trust in AI system) [91], * Safety Incidents (number of safety-related issues) [92], * Error Reduction Rate (decrease in errors due to AI assistance) [4], * Confidence (level of user trust in AI recommendations) [93] Task Allocation Complementarity AI handles data analysis and uncertainty estimation; Humans provide expertise and clarification. * Query Efficiency (amount of information gained per query), * Expertise Utilization (degree to which human expertise is leveraged) [84] Flexibility Adjusting query strategy based on interaction dynamics. * Adaptability Score (same as in Interaction), * Dynamic Task Allocation (ability to adjust task distribution based on real-time needs) [3] Efficiency Streamlining the learning process through targeted queries. * Model Improvement Rate (rate of improvement in AI performance) [65, 84], * Learning Curve (same as in Goals), * Resource Utilization (efficiency in using computational resources) [65] Responsiveness AI’s ability to rapidly adjust to new inputs and feedback. * Response Time (time taken to respond to queries or feedback), * Real-time Performance (ability to handle time-sensitive tasks) [82, 94] Collaborative Decision Making Joint decisions on data labeling and correction of AI’s model. * Impact of Corrections (measured improvement in system performance through human intervention) [95, 96], * Decision Effectiveness (impact of decisions on overall task success) [97, 98] Continuous Learning AI’s ability to learn from human input and adjust its model. * Model Improvement Rate (same as in Efficiency), * Learning Curve (same as in Goals), * Knowledge Retention (ability to retain learned information over time) [99] Mutual Support AI’s ability to provide support to humans in the learning process; Humans support by providing information. * Error Reduction Rate (decrease in errors due to AI assistance) [4], * Confidence (level of user trust in AI recommendations) [100], * Task Completion Time (with and without AI support) [101] Robustness Ensuring AI system’s robustness and reliability. * Adversarial Robustness (resilience to adversarial attacks) [102], * System Reliability (overall system uptime and fault tolerance) [103]
Table 2: Factors and Subfactors of Human-AI Collaboration with Associated Metrics for Evaluation
Goals
The Goals factor ensures that the collaboration has a clear direction and that both human and AI efforts are aligned towards shared outcomes. This factor encompasses two key subfactors:
  • Individual Goals: The specific aims of each participant. For the AI, this could be learning efficiently or maximizing accuracy. For humans, it might involve providing effective teaching or achieving task objectives. Misalignment of individual goals can hinder collaboration, as highlighted in studies on migrating to HAIC, which emphasize the importance of goal congruence [40].
    • Learning Curve: Measures the rate of improvement over time, indicating the effectiveness of human teaching and AI learning. This is plotted by graphing the performance of the AI (like accuracy) against the number of training samples or training iterations. A steep learning curve suggests the AI is learning effectively [81].
    • Prediction Accuracy: The proportion of correct predictions made by the AI model out of all predictions. This can be measured using metrics like accuracy, F1 score, or area under the ROC curve, depending on the nature of the task [86].
    • Teaching Efficiency: The gain in AI performance per unit of teaching effort by the human. This could be measured by tracking improvements in the AI’s accuracy or error rate after receiving human input. This concept is explored in works like that of Cakmak and Thomaz [85]
  • Collective Goals: Shared objectives that maximize the overall performance of the HAIC system. For example, in a medical diagnosis system, the collective goal might be to improve patient outcomes through accurate and timely diagnosis. These collective goals are often broader than individual goals and necessitate the combined efforts of both humans and AI.
    • Overall System Accuracy: It can be calculated as a weighted average of different performance metrics, reflecting the relative importance of each task in achieving the overall goal. Evaluates the general accuracy of the AI-human collaborative system, tailored to the specific objectives of the system. [65].
    • Objective Fulfillment Rate: The percentage of predefined goals that are successfully achieved within the HAIC collaboration. This metric is often used in project management contexts to track progress towards overall objectives and can be adapted to assess the success of HAIC collaborations [87].
Schneider et al. [97] provide empirical evidence on quality goal rankings, while Wang et al. [40] highlight the necessity of goal alignment in collaborative environments. Fiebrink et al. [104] illustrate how human input influences AI’s learning processes, emphasizing the mutual dependence of individual and collective goals.
Interaction
The Interaction factor underscores the critical communication mechanisms through which humans and AI systems exchange information, provide mutual feedback, and adaptively respond to each other’s inputs. Key subfactors within the Interaction factor include:
  • Communication Methods: This involves evaluating the clarity, accessibility, and intuitiveness of the interfaces and protocols through which humans and AI interact. For instance, the clarity of communication in a medical diagnosis system might be assessed by how well the AI can explain its reasoning to the doctor in an understandable way [105, 90]. Effective communication methods ensure efficient information exchange and mutual understanding, which can be evaluated using metrics such as:
    • Clarity of Communication: Can be assessed through qualitative methods like expert review or user feedback, or through quantitative measures like the proportion of correctly interpreted queries or the time taken to resolve ambiguities.
    • Ease of Use: Often measured through Likert scale questionnaires or user interviews to gauge the user’s subjective experience with the system.
  • Feedback Mechanisms: These mechanisms assess how input from human operators is used to guide AI learning and behavior adjustments. For instance, in a customer service chatbot, feedback mechanisms could involve analyzing how user ratings or corrections influence the chatbot’s future responses. It’s crucial that AI systems can not only receive but also meaningfully integrate feedback to enhance their interactions with humans [83, 88]. The effectiveness of feedback mechanisms can be measured through:
    • Frequency of Feedback (number of feedback instances): This is a simple count of how often users provide feedback, indicating engagement and potentiallearning opportunities for the AI, as highlighted in Amershi et al.’s (2014) [83] work on power to the people.
    • Feedback Quality (relevance and specificity of feedback): This can be measured qualitatively through expert review or quantitatively by analyzing the correlation between feedback and subsequent AI performance improvements as suggested by Dautenhahn (2007) [88] in the context of socially intelligent robots.
    • Feedback Impact (measured change in AI performance due to feedback): This quantifies the tangible effect of feedback on AI behavior or output quality, often measured by improvements in accuracy, precision, or other relevant metrics.
  • Adaptability: This measures the AI’s ability to modify its behaviors based on feedback and changes in the environment or task conditions. Adaptability is essential for AI systems to remain effective as collaboration needs evolve [106, 107]. For example, in an autonomous vehicle, the AI system should be able to adapt its driving style based on the driver’s feedback or changes in traffic conditions. Adaptability can be evaluated through:
    • Adaptability Score (composite metric of flexibility and responsiveness): This metric combines measures of how quickly and effectively the AI responds to changes and how flexibly it adjusts its strategies, which can be derived from studies like that of El-Assady et al. (2022) [90] on communication biases in HAIC.
    • Strategy Adaptation Effectiveness (improvement in query selection based on feedback): In systems that involve querying for information, this metric measures whether the AI learns to ask more effective questions based on human responses, potentially incorporating insights from research on semantic interaction challenges by Wenskovitch et al. (2020) [89].
  • Trust and Safety: Ensuring trust in the AI system and the safety of users is critical for effective HAIC. Metrics to evaluate trust and safety include:
    • Trust Score: The level of user trust in the AI system, often measured through surveys and trust scales [100].
    • Safety Incidents: The number of safety-related issues encountered, used to assess the reliability and safety of the AI system.
Ethical considerations such as bias, fairness, and transparency are critical to the success of HAIC. Significant research has underscored the complexity of achieving fair and unbiased AI systems. For example, Barocas et al. [108] discuss the ethical implications of bias in ML systems, highlighting the need for transparent and accountable AI practices. Holstein et al. [109] emphasize the importance of fairness in AI design and implementation, advocating for continuous monitoring and evaluation to mitigate bias.
Task Allocation
Task allocation in HAIC focuses on strategically distributing responsibilities between humans and AI to leverage the strengths of each party and optimize overall performance. Effective task allocation enhances efficiency, minimizes errors, and increases user satisfaction, making it a critical component in the success of HAIC systems [110]. Key subfactors in Task Allocation, and its relevant metrics include:
  • Complementarity: Aligning tasks with the unique capabilities of humans and AI. For instance, AI’s analytical strengths can complement human decision making and creativity [111].
    • Query Efficiency: Indicates the number of queries the AI needs to reach a certain level of accuracy, with fewer queries signifying more efficient learning [84].
    • Expertise Utilization: The degree to which human expertise is leveraged, often measured qualitatively through observation or interviews, or quantitatively through metrics like the percentage of decisions where human input is decisive [26].
  • Flexibility: The ability to adjust task allocation dynamically in response to changing circumstances or new challenges is crucial for maintaining operational effectiveness [112]. Research by Dubey et al. (2020) [3] explores how AI can be used to optimize task allocation for improved efficiency.
    • Dynamic Task Allocation: The ability to adjust task distribution based on real-time needs, assessed by how quickly and effectively the system adapts to changes in workload, priorities, or available resources, as explored in studies on human-robot collaboration by Tsarouchi et al. (2017) [112].
  • Efficiency: Focuses on maximizing productivity and resource utilization by optimizing task distribution, often allowing AI to automate routine tasks [111, 39]. Efficiency can be measured by:
    • Model Improvement Rate: The rate of improvement in AI performance, calculated by taking the difference between the AI’s performance at a later time point and its performance at an earlier time point. This difference is then divided by the time interval between these two points. This can be applied to various AI performance metrics like accuracy or error rates, and its significance in streamlining HAIC has been explored by studies such as Vossing et al. (2022) [65].
    • Resource Utilization: The efficiency in using computational resources like CPU time, memory, or network bandwidth. This can be measured as a percentage of utilized resources compared to the total available, and can be a key indicator of overall system efficiency.
    • Learning Curve: The rate of improvement in AI performance over time, as defined earlier in the Goals section.
  • Responsiveness: The capability of systems to rapidly adjust to new requirements or unexpected situations is essential for maintaining workflow continuity, as discussed by Johannsmeier and Haddadin (2016) [113]. Responsiveness can be evaluated through:
    • Response Time: The average time taken for the AI or human to respond to queries or feedback. A lower response time indicates quicker and more effective communication, as highlighted in studies on human-robot interaction by Hemmer et al. (2022) [111, 82].
    • Real-time Performance: The ability to handle time-sensitive tasks, measured by factors like throughput (tasks completed per unit time) and latency (delay in response) [94]. This is particularly important in applications like real-time decision support systems and autonomous vehicles.
    • Learning Curve: The rate of improvement in AI performance over time, as defined earlier in the Goals section.
  • Collaborative Decision Making [111]: Involves integrating the insights and expertise of both humans and AI to make informed decisions, ensuring that all parties contribute to problem- solving.
    • Impact of Corrections: Measured improvement in system performance resulting from human intervention, often quantified by changes in accuracy or error rates after corrections are made. This can be used to assess the effectiveness of human input in refining AI outputs [96, 95].
    • Decision Effectiveness: The impact of decisions on the overall task success, measured by comparing outcomes of decisions made with and without AI collaboration. This metric can be customized based on the specific goals of the HAIC system [98, 97].
  • Continuous Learning: AI systems should evolve by learning from past interactions to enhance future performance and adaptation. Yin et al. (2021) highlight the importance of continuous learning for optimizing task allocation in dynamic environments [114].
    • Knowledge Retention: The AI’s ability to retain learned information over time, often assessed by testing the AI’s performance on tasks after a period of non-use. This metric is important for understanding the long-term impact of human feedback and training on the AI system.
    • Model Improvement Rate: (same as in Efficiency)
    • Learning Curve: (same as in Goals)
  • Mutual Support: Promoting a supportive environment where both humans and AI can offer reciprocalassistance enhances the overall collaboration effectiveness [111]. Mutual support can be evaluated through:
    • Error Reduction Rate: Measures the decrease in errors or mistakes made by the AI following human intervention, showing the effectiveness of human corrections  [4].
    • Confidence: The level of user trust in AI recommendations, often measured using Likert-scale questionnaires or surveys [100]. Higher confidence scores indicate a stronger belief in the AI’s capabilities and a greater likelihood of accepting its recommendations.
    • Task Completion Time: Comparison of task completion times with and without AI support [101]. This metric helps assess the efficiency gains resulting from HAIC.
  • Robustness: Ensuring the AI system performs reliably under various conditions, including adversarial attacks and different domains, is crucial for HAIC effectiveness. Robustness can be evaluated through:
    • Adversarial Robustness: The resilience of the AI system to adversarial attacks, measured through testing against adversarial examples.
    • System Reliability: The overall system uptime and fault tolerance, measured through system logs and reliability tests.
This structured approach, detailing subfactors and metrics, ensures a comprehensive evaluation of HAIC systems, enabling continuous improvement and effective collaboration between humans and AI.
The Table 3 offers definitions and calculations for each metric described in Table 2, facilitating a full knowledge of the assessment metrics used in the structured framework for Human-AI Collaboration (HAIC). These metrics are essential for quantitatively evaluating the efficacy, efficiency, and adaptability of HAIC systems across multiple domains.
{adjustwidth}
[…]
Table 3: Evaluation Metrics for Human-AI Collaboration
4.3 Building the Framework & Quantifying the Evaluation of HAIC
Building upon influential research by Fischer [115], Wenskovitch and North [89], and Sharma et. al [49], this framework offers a structured approach to Human-AI Collaboration (HAIC) evaluation. This is achieved using a decision tree (Fig. 2), that guides users through a series of questions to identify the most relevant metrics for their specific HAIC system. This combines the factors, subfactors, and metrics identified in the previous sections, based on the HAIC modes described earlier in Section 2.3, i.e., AI-Centric, Human-Centric, and Symbiotic.
Figure 2: Decision Tree Framework for Evaluating Human-AI Collaboration
Notice in Fig. 2 that each metric has a distinct color based on the main factor that is coming from the Table 2. This color-coding system helps users quickly identify the primary factor each metric belongs to, facilitating a more intuitive and user-friendly evaluation process. The framework is designed to be adaptable and scalable, accommodating a wide range of HAIC systems and applications. By providing a structured and comprehensive evaluation approach, this framework aims to enhance the understanding and assessment of HAIC, fostering the development of more effective and collaborative human-AI systems.
AI-Centric Collaboration
Questions in this mode emphasize the AI’s autonomous capabilities and performance. We focus on efficiency metrics (like Model Improvement Rate) and uncertainty handling, as these are core challenges for AI systems operating primarily independently. The AI-Centric mode is particularly relevant for systems where AI plays a dominant role in decision-making, such as predictive analytics and automated systems.
How efficiently does the AI make decisions independently?
  • Learning Curve: Measures the rate of improvement in AI performance over time.
  • Model Improvement Rate: Measures improvement over time, indicating decision-making efficiency.
  • Query Efficiency: Evaluates how much information is gained with minimal queries.
  • Resource Utilization: Assesses how well the AI manages computing power.
  • Response Time: Tracks how quickly the AI makes decisions.
  • Real-time Performance: Evaluates the AI’s ability to handle time-sensitive tasks.
  • Objective Fulfillment Rate: Measures the percentage of tasks successfully completed.
How well does the AI handle uncertainty?
  • System Accuracy: Measures overall performance, partially reflecting uncertainty handling.
  • Error Reduction Rate: Indicates if the AI makes fewer mistakes over time.
  • Prediction Accuracy: Quantifies correct outputs specifically in uncertain scenarios.
  • Feedback Impact: Shows how much human guidance improves AI decisions under uncertainty.
Human-Centric Collaboration
Here, the questions center on how well the AI communicates with and supports the human user. Metrics like Clarity of Communication and Ease of Use are essential for a positive user experience. The Human-Centric mode is particularly relevant for systems where human input is crucial for decision-making, such as in creative fields or complex problem-solving.
How clear and usable is the AI’s communication?
  • Clarity of Communication: Directly assesses if queries and responses are understandable.
  • Ease of Use: Measures overall user satisfaction with the interaction.
  • User Satisfaction: Indicates if the AI’s communication methods are acceptable to humans.
  • Frequency of Feedback: Quantifies the number of feedback instances.
How effectively does the AI support human tasks?
  • Teaching Efficiency: Gauges how much effort is needed to train the AI.
  • Error Reduction Rate: Quantifies how much the AI helps in reducing human errors.
  • Confidence: Measures user trust in the AI’s recommendations.
  • Task Completion Time: Indicates if AI assistance makes tasks faster.
Symbiotic Collaboration
This mode’s questions strike a balance between adaptability and task allocation. We ask about the system’s flexibility in response to feedback (Adaptability Score) and the effectiveness of resource distribution, reflecting the collaborative core of Symbiosis.
How easily does the AI adapt to human feedback?
  • Adaptability Score: A composite metric reflecting flexibility and responsiveness.
  • Strategy Adaptation Effectiveness: Measures how well the AI changes approaches based on input.
  • Feedback Quality: Assesses how valuable the provided human feedback is.
  • Feedback Impact: Shows the improvement resulting from human guidance.
How well are tasks allocated to optimize both human and AI strengths?
  • Dynamic Task Allocation: Reflects the ability to shift task distribution in real-time.
  • Query Efficiency: Assesses how well information is requested to leverage human expertise.
  • Expertise Utilization: Measures the degree to which human skills are leveraged.
  • Knowledge Retention: Evaluates if the AI learns effectively from human input.
  • Model Improvement Rate: Measures how much human input contributes to AI progress.
  • Decision Effectiveness: Indicates the impact of decisions on overall task success.
Advantages of the Decision Tree Approach
This methodology offers several benefits:
  • User-Friendly: Simplifies the evaluation process, making it accessible to users with varying levels of expertise.
  • Tailored Assessment: Ensures that the evaluation is relevant to the specific characteristics and goals of the HAIC system.
  • Adaptable Framework: Accommodates the evolving nature of HAIC technologies and practices.
Although the decision-tree based framework offers a solid foundation for evaluating HAIC, the field continues to make significant advancements. As AI systems advance and expand their range of uses, various challenges and prospects arise. In order to maintain effectiveness, it is necessary for our evaluation methods to also adjust and change.
Weighting Mechanism & Overall Score Calculation
While the decision tree and associated metrics lead the evaluation process, we recognize the importance of a weighting mechanism to provide a more comprehensive evaluation of HAIC effectiveness. This system would weight each metric according to its relevance to the specific HAIC mode and the evaluation’s criteria. For example, in an AI-Centric system, accuracy might be assigned a higher weight than user satisfaction. Conversely, in a Human-Centric system, ease of use and trust could be prioritized. In the Symbiotic mode, the weights might be distributed more evenly across metrics reflecting collaboration and mutual benefit.
The overall score for a HAIC system could then be calculated as the weighted average of the individual metric scores. However, determining appropriate weights requires further research and validation across different real - world paradigms.
Several studies in the literature have utilized weighting mechanisms to evaluate complex systems. For example, Kamar et al. (2012) [116] employed a weighted approach to combine human and machine contributions in a crowdsourcing system, emphasizing the importance of weighting in achieving balanced evaluations. Similarly, Amershi et al. (2014) [83] discussed the role of weighting different interaction metrics to assess the effectiveness of human-AI collaboration tools, highlighting how different priorities can shift the weighting schema.
Moreover, Vossing et al. (2022) [65] implemented a weighting mechanism to balance transparency and performance in AI systems, demonstrating that context-specific weighting can enhance the evaluation’s relevance and accuracy. These examples underscore the practical value of weighting mechanisms and support their inclusion in our proposed framework.
By incorporating a validated weighting mechanism, our framework ensures a nuanced and context-sensitive assessment of HAIC systems, aligning with best practices in the literature. Future research should focus on refining these weights through empirical studies across various HAIC applications to enhance the robustness and applicability of the evaluation framework.
5 Implications of the HAIC Evaluation Framework to Real-World Domains
5.1 Manufacturing
The manufacturing industry’s focus on safety, accuracy, and productivity closely corresponds with the Symbiotic mode of HAIC, which requires the collaboration of human expertise and AI skills to get the best results. This mode prioritizes metrics like Adaptability Score and Confidence. Recent study highlights a shift from focusing primarily on productivity- centered measures to a more worker-centric approach, placing greater emphasis on safety and trust. Research such as Rabby et al.’s (2020) [117] time-driven performance-aware trust model and Okamura and Yamada’s (2020)  [118] adaptive trust calibration approach exemplify this change. These works emphasize the importance of the Interaction factor, specifically focusing on the subfactors of Communication Methods and Adaptability. They achieve this by creating systems that can adapt to human actions and uncertainties. The primary emphasis of Amin et al.’s (2020)  [119] mixed-perception method is on safety, which aligns with the Task Allocation factor’s concentration on Complementarity and Mutual Support. This strategy involves the collaboration between human and AI systems to promote worker protection. The focus placed by Ming Hou (2021)  [120] on context-based interface design provides additional support for the Symbiotic method. Our methodology proposes the following essential criteria for evaluating HAIC systems in manufacturing: Adaptability Score, Error Reduction Rate, Confidence, and Task Completion Time. These measurements will assess the systems’ capacity to adapt to human operators, enhance safety, foster worker confidence, and ultimately optimize industrial processes.
5.2 Healthcare
The integration of AI in healthcare, particularly in medical imaging, often operates in a collaborative mode, where AI augments human expertise to achieve more accurate diagnoses. Although AI is highly effective in medical image classification and segmentation  [121], human skill is still essential for the interpretation and ultimate diagnosis. Contributions to this collaborative approach are made by studies such as Zheng et al.’s (2023)) [122], Eldin & Kaboudan’s (2023) [123] study of AI-driven platforms, and Aleid et al.’s (2023) [53] work on tumor identification. These developments specifically focus on the Goals aspect of the HAIC framework, with a special emphasis on enhancing diagnostic Accuracy and Efficiency. The Communication Methods subfactor is crucial as AI systems must provide unambiguous and comprehensible outputs, such as the Clarity of Communication. In the ICU setting, as explored by Dziorny et al. (2022) [124], the quality of Decision Support provided by AI becomes crucial, emphasizing the need for evaluating metrics like Decision Effectiveness. Our approach emphasizes key parameters such as System Accuracy, Prediction Accuracy, Response Time, and Clarity of Communication to successfully evaluate these collaborative HAIC systems. Future study should investigate the impact of such systems on the workflow of radiologists and oncologists, specifically in relation to Decision Support and Time-to-Diagnosis. Furthermore, conducting research to examine any possible effects on Patient Communication would be important.
5.3 Finance
The finance sector largely employs HAIC in a Symbiotic manner, leveraging AI’s analytical capabilities alongside human financial experience. This collaboration is evident in the work of Atadoga et al. (2024) [125], Al-Fatlawi et al. (2024)  [54], Jain (2023) [126], and Ali et al. (2022) [127]. These studies highlight the importance of Goals such as Accuracy and Efficiency in minimizing fraud and operational costs. However, the Symbiotic nature of this collaboration also emphasizes the Interaction factor, specifically the subfactor of Clarity of Communication, as AI-generated insights must be clearly communicated to human decision-makers.The Task Allocation factor is evident in the division of labor, with AI systems often identifying potential fraud while human experts analyze complex cases. The Error Reduction Rate is a key metric to assess the effectiveness of AI in this context. Future research should investigate the ethical and security implications of these systems, along with their long-term impacts on client relationships and the quality of financial decisions.
5.4 Education
HAIC in Education demonstrates a hybrid of Symbiotic and Human-Centric paradigms, with AI playing a supportive role in enhancing both teaching practices and learner experiences. Studies by Ifenthaler and Schumacher (2023) [128], Ji et al. (2023) [129], Wang et al. (2020)  [40], and Sharma et al. (2023) [49] highlight these modes. These works target Goals such as enhancing Teaching Efficiency and achieving personalized learning through Individualized Goals. The Interaction factor is crucial, as effective communication between AI tutors and learners impacts the learning process, necessitating assessment of Clarity of Communication and Ease of Use. The Lumilo project, investigated by Holstein and Aleven (2022) [12], shows Symbiotic collaboration for improved learning outcomes, where real-time analytics provided by AI enhance the teacher’s ability to provide Mutual Support to students. Evaluating the impact of these systems requires metrics like Task Completion Time and Learning Curve to ensure a comprehensive understanding of both student and teacher experiences. Further research should emphasize fostering student agency and critical thinking within HAIC frameworks, moving beyond passive consumption of AI-generated tools toward active engagement and co-creation of knowledge, as advocated by Fadiya (2023)  [130].
6 The Unique Challenge of Evaluating Creative and Linguistic AI
As seen in the previous sections, establishing comprehensive frameworks for assessing HAIC is vital, but certain applications pose distinct challenges due to their unique nature. Two prominent examples are Large Language Models (LLMs) and Generative AI in the Arts, where AI’s impact on human creativity, expression, and communication requires specialized evaluation methodologies.
While our existing framework provides a foundation for HAIC assessment, these creative and linguistic domains raise questions about the adequacy of traditional metrics. Can quantitative measures, such as the ones listed in Table 3 fully capture the nuances of artistic collaboration or the impact of AI on human language? How can we evaluate the ethical implications of AI- generated art or the potential for bias in LLMs? These challenges necessitate a more nuanced and tailored approach to evaluation, as we will explore in the following sections.
6.1 Large Language Models (LLMs)
Large Language Models (LLMs) have revolutionized text generation, translation, and dialogue. The GPT-3 [131] family, including ChatGPT and GPT-4, showcases the power of these models to engage in complex natural language processing [132]. While their potential for transforming Human-AI Collaboration (HAIC) is vast, evaluating these interactions requires a critical lens and the development of new methodologies [133].
Ensuring the interpretability and fairness of LLMs is crucial. Binz & Schulz’s (2022) [134] study reveals that even when LLMs perform well on cognitive tasks, they might employ reasoning patterns that are opaque or reflect biases embedded in their training data. Evaluation frameworks must incorporate methods for assessing how well LLMs explain their outputs and whether biases are mitigated over time. Additionally, the quality of human interaction with LLMs is paramount. Using GPT-3 to simulate human interactions, as in the work by Aher, Arriaga, & Kalai (2022) [56], demonstrates the value of testing how well LLMs adapt to different users and communication styles.
LLMs promise to augment cognitive tasks, enhance creativity, and personalize learning. Webb, Holyoak, & Lu (2022) [135] demonstrate an LLM’s capability for analogical reasoning, a core element of problem-solving. Further research should evaluate how such abilities impact HAIC success in real-world tasks. Regarding creativity, Douglas (2023) [136] highlights LLMs’ transformative role, inviting us to develop metrics for assessing how these tools amplify human ideation and artistic expression. In education, Liu et al. (2023) [137] explore the potential of ChatGPT and GPT-4 models to tailor learning experiences. Evaluating their effectiveness in this context will necessitate student-centered metrics focused on knowledge gains and learner autonomy.
Longitudinal studies are vital for a nuanced understanding of LLM-based HAIC. Work by Yin et al. (2023) [138] on LLM self-knowledge suggests ways to track how both the AI’s capabilities and the user’s trust in the system evolve over extended periods of interaction. The work of Dwivedi et al. (2023) [139] highlights crucial challenges posed by ChatGPT, reminding us that even with the potential benefits, addressing societal implications and ethical use is paramount.
The development of robust evaluation frameworks for LLM-powered HAIC is a rapidly evolving field. By addressing interpretability, bias, the quality of interaction, and the transformative potential of LLMs on human cognition, such frameworks will drive the creation of responsible, effective, and truly collaborative AI systems.
6.2 Generative AI in the Arts
Generative AI is revolutionizing artistic expression by enabling algorithms to generate artworks in various modalities, including visual art, music, and poetry. This technology transforms AI systems from tools into active co-creators, expanding the boundaries of what artists can achieve. Notable examples include AI-generated paintings, musical compositions, and even interactive installations tailored to audience responses. Studies like that of Gmeiner, Holstein, & Martelaro (2022) [140] suggest that understanding how effective human teams learn can provide a framework for designing generative AI systems that maximize collaborative potential.
Evaluating HAIC in the context of generative AI for artists poses several challenges. Traditional art metrics such as aesthetic quality, originality, technical skill, audience reception, market value, and art historical significance provide a starting point for this evaluation [141]. These metrics, which typically focus on the final artwork, can be adapted to also consider the collaborative process between the artist and the AI. Verheijden & Funk’s (2023) [57] work highlights the importance of assessing the generative AI’s impact on the design process itself - how it boosts inspiration, facilitates communication, and shapes the creative journey.
The field necessitates methodologies that evaluate AI-assisted artistic creation holistically. Frameworks inspired by HAIC models developed for other domains will require adaptation to consider factors unique to artistic collaboration. He, Shrestha, Puranam, & Miron-Spektor (2023) [142] propose a novel theory of co-creativity as a joint search process, offering an insightful lens for analyzing the dynamic between human artist and generative AI system.
The rise of AI-assisted art brings both ethical considerations and profound societal implications. Issues of copyright, ownership, and the meaning of authenticity in this new landscape demand attention. Epstein et al. (2023) [143] call for interdisciplinary investigation into generative AI’s cultural, economic, and legal impacts. Additionally, as AI tools become ubiquitous, reevaluating the definition of "artist" and determining the skills that foster success in this evolving domain become essential considerations.
As generative AI in art continues to mature, its transformative potential is immense. We might witness a shift towards real-time, personalized art experiences, requiring HAIC evaluation models to be responsive and adaptive. Suh, Youngblom, Terry, & Cai (2021) [144] offer intriguing insights into how AI might reshape social interactions within creative settings. Moreover, focusing on documenting the artist-AI creative process itself could become as paramount as evaluating any final product.
7 Limitations in Investigating Behavioral Factors
Our study provides a thorough framework to evaluate Human-AI Collaboration (HAIC) in different sectors, emphasizing goals, interaction dynamics, and task allocation. However, it does not explicitly examine behavioral aspects as a distinct main category of analysis. The exclusion of behavioral aspects of HAIC is not because they are unimportant but is due to methodological and scope concerns.
The study of behavioral aspects in the context of HAIC involves a multidisciplinary approach, incorporating knowledge from psychology, sociology, cognitive science, and information technology to address its complexity. The complexity of the research requires interdisciplinary collaboration and skills outside the current study’s context.
Behavioral aspects are important but can vary significantly and depend greatly on the context. The wide range of human behavior, shaped by cultural, psychological, and situational elements, presents a difficulty in establishing a uniform or universal evaluation framework for this topic. Extensive empirical study, including longitudinal studies and ethnographic methodologies, like those in Yang, Carroll, & Dragan (2022) [55], would provide a richer picture of how AI adaptability impacts human trust over time and how trust, in turn, shapes collaborative goals.
Our focus on the structural elements of HAIC, like goals, interaction, and task allocation, enables a more objective examination based on observable and quantifiable results. Behavioral elements demand methodological techniques and analytical frameworks different from those used in our study due to their subjective interpretations and qualitative judgments.
Behavioral variables are recognized as crucial for the success and efficacy of HAIC systems. We support more study focused on understanding these factors thoroughly. These studies could provide useful insights into the mutually beneficial connection between human views, beliefs, and behaviors and AI collaboration. This would help in creating more sophisticated, efficient, and human-centered HAIC systems.
Behavioral aspects are crucial in understanding and optimizing HAIC. However, they were intentionally excluded as a primary emphasis in this study to keep the scope focused and manageable. We urge further study to fill this gap, expanding on the groundwork established by our approach to investigate the complex interaction between human behavior and AI collaboration.
8 Conclusion
This paper has explored the current state in the evaluation of HAIC and has illuminated a notable gap in the current HAIC evaluation methodologies: the absence of a framework that effectively integrates qualitative metrics alongside quantitative ones, applicable across a wide range of sectors. Our investigation revealed that existing evaluation practices are predominantly domain-specific, lacking the structured adaptability required for comprehensive HAIC assessment. The framework proposed herein endeavors to address this void by offering a structured approach to HAIC evaluation, one that is predicated on the mode of collaboration rather than the sector, application, or AI method used. This decomposition of evaluation into smaller, manageable components presents a novel way to scrutinize HAIC in its multifaceted nature.
The significance of our framework lies in its universal applicability, affording a comparative lens to examine different collaboration methods. By transcending domain-specific limitations, it paves the way for a standardized evaluation approach that could be adopted across various sectors, enhancing the consistency and reliability of HAIC assessments.
While the framework stands as a theoretical construct, its real-world implications remain untested. The future trajectory of our research is clear: to implement this framework as a software platform and apply it across diverse sectors and collaboration modes to validate its efficacy and adaptability. Such empirical endeavors will be crucial in refining the framework and ensuring its robustness.
Our study acknowledges its limitations, particularly the omission of behavioral factors and ethical considerations from the framework. These elements, while outside our current scope, are critical dimensions of HAIC that warrant attention in future research.
As we look to the horizon, our belief is steadfast in the potential of this framework to significantly advance the methods of evaluating HAIC. As this interdisciplinary field evolves, a standardized framework becomes increasingly crucial for its adoption from industry to society. We anticipate that our contribution will serve as a catalyst for future studies, spurring the development of HAIC evaluations that are as dynamic and intelligent as the collaborations they aim to measure.
Acknowledgments
The work leading to these results has received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101120218, project HumAIne (Hybrid Human-AI Decision Support for Enhanced Human Empowerment in Dynamic Situations). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@
 Abstract
-The use of artificial intelligence (AI) in working environments with individuals, known as Human-AI Collaboration (HAIC), has become essential in a variety of domains, boosting decision-making, efficiency, and innovation.
-Despite HAIC’s wide potential, evaluating its effectiveness remains challenging due to the complex interaction of components involved.
-This paper provides a detailed analysis of existing HAIC evaluation approaches and develops a fresh paradigm for more effectively evaluating these systems.
-Our framework includes a structured decision tree which assists to select relevant metrics based on distinct HAIC modes (AI-Centric, Human-Centric, and Symbiotic).
-By including both quantitative and qualitative metrics, the framework seeks to represent HAIC’s dynamic and reciprocal nature, enabling the assessment of its impact and success.
-This framework’s practicality can be examined by its application in an array of domains, including manufacturing, healthcare, finance, and education, each of which has unique challenges and requirements.
-Our hope is that this study will facilitate further research on the systematic evaluation of HAIC in real-world applications.
-Keywords Human-AI Collaboration (HAIC)  $\cdot$ HAIC Evaluation  $\cdot$ Symbiotic AI Systems  $\cdot$ AI Metrics  $\cdot$ AI Evaluation Framework
+The use of artificial intelligence (AI) in collaborative working environments, termed Human-AI Collaboration (HAIC), has become vital across diverse domains, enhancing decision-making, efficiency, creativity, and innovation.
+Despite its wide potential, evaluating HAIC effectiveness poses significant challenges due to the complex, dynamic, and reciprocal interactions between humans and AI systems.
+This paper provides a comprehensive analysis of existing HAIC evaluation methods and introduces a novel, domain-agnostic framework for systematically assessing these systems.
+To support structured assessment, our framework incorporates a decision-tree-based approach that helps select relevant metrics tailored to distinct HAIC modes (AI-Centric, Human-Centric, and Symbiotic).
+By integrating both quantitative and qualitative evaluation criteria, the framework aims to address key gaps in HAIC assessment methodologies.
+We highlight the unique challenges posed by evaluating creative and linguistic AI applications, such as large language models and generative AI in the arts, underscoring the need for tailored evaluation approaches in these emerging areas.
+This work lays the foundation for future research and empirical validation, offering a structured methodology to enhance the evaluation of HAIC systems.
+Keywords Collaboration  $\cdot$ Evaluation  $\cdot$ Symbiosis  $\cdot$ Metrics  $\cdot$ Artificial Intelligence  $\cdot$ Interaction  $\cdot$ Adaptability
 1 Introduction
-The fast advancement of Artificial Intelligence (AI) has fundamentally transformed several domains, blurring the distinctions between human and machine roles and giving rise to new interactions [1].
-The notion of Human-AI Collaboration (HAIC) has emerged as a pivotal framework, as evidenced by the research conducted by [2] and [3].
-These studies identify that HAIC has significant potential for enhancing decision-making, increasing efficiency, and fostering innovation.
-They therefore aim at fostering and facilitating effective cooperation between people and AI systems to achieve common goals.
-On the other hand, assessing the benefits of this cooperation presents its own challenges, owing to the multiple complex factors involved.
-Traditional Human-Machine evaluation frameworks have primarily focused on quantifiable metrics such as task performance, response time, and user satisfaction within human-machine systems [4, 5].
-These frameworks typically view the machine as a tool and assess how effectively humans can control or interact with this tool to achieve desired outcomes.
-Human-AI Collaboration (HAIC) represents a significant change in which humans and AI work together in a more integrated and mutually beneficial manner.
-HAIC focuses on more than just human dominance over machines; it emphasizes a collaborative partnership where individuals and AI systems actively contribute to shared results.
-The significance of this partnership has been highlighted in several domains, such as manufacturing [6, 7], financial services [8, 9], healthcare [10, 11], education [12], and others, demonstrating notable advancements in both technological innovation and practical implementation [13, 14].
-HAIC is defined by the process of making decisions together, learning from each other, and adapting, which requires a more comprehensive and nuanced evaluation approach [15, 16].
-To effectively evaluate HAIC, it is crucial to consider not only the traditional metrics of task success but also the quality of interactions within the team, the processes by which decisions are made, and crucial ethical aspects such as the transparency of algorithms and the mitigation of biases [17].
-This involves assessing how AI systems influence human decisions, how humans and AI adapt to each other’s capabilities, and how ethical considerations are integrated into collaborative processes.
-In this paper, we aim at addressing these evolving needs by critically analyzing current evaluation methods in human-AI interaction and proposing a new framework designed to capture the dynamic and reciprocal nature of human-AI collaborations.
-This framework seeks to recognize and measure both the individual and joint contributions of humans and AI to collaborative outcomes, thus providing a more comprehensive understanding of HAIC effectiveness.
-The contributions of this paper are therefore the following.
-  • Overview of the Literature: Extending the foundational work of Madni [2] and Dubey et al.
-[3], we provide a review of HAIC evaluation methods, identifying gaps and research opportunities in the existing literature.
-  • Novel Evaluation Framework: Inspired by the pioneering frameworks of Rouse et al.[18] and Parasuraman et al.[19, 4], we propose a new, adaptable evaluation framework for HAIC that incorporates both quantitative and qualitative measures, emphasizing ethical considerations and providing a flexible structure for diverse HAIC applications.
-  • Domain-Specific Insights: Through detailed examination of applications in healthcare, finance, and education, as informed by the studies of Tschandl et al.[10], Buckley and Lui[8, 9], and Holstein et al.
-[12], we demonstrate the applicability and customization of our framework to meet domain-specific needs.
-Structure of the Paper
-Following this introduction, Section 2 discusses the fundamentals of Human-AI Collaboration, including various modes of interaction.
-Section 3 reviews the existing literature on HAIC, highlighting advancements in evaluation frameworks and tools, as well as the diversity and challenges in domain-specific evaluations.
-Section 4 presents a structured framework for evaluating HAIC, detailing factors, subfactors, and metrics for assessment.
-Section 5 applies this evaluation framework to real-world domains such as manufacturing, healthcare, finance, and education.
-Section 6 addresses the unique challenges of evaluating creative and linguistic AI, including large language models and generative AI in the arts.
-Section 7 discusses the limitations in investigating behavioral factors.
-Finally, Section 8 discusses the implications of the HAIC evaluation framework for future research and practice.
-2 Understanding the Fundamentals of Human-AI Collaboration
+Evaluating Human-AI Collaboration (HAIC) remains a significant challenge due to the complex interactions between humans and AI systems.
+The effectiveness of HAIC depends not only on system performance but also on the quality of human-AI interaction, trust, and adaptability.
+However, existing evaluation methods often fail to capture these multidimensional aspects, making it difficult to establish comprehensive assessment criteria.
+To address this issue, a structured framework that provides a foundation for systematically assessing HAIC across various domains, is introduced in the paper.
+Is is inspired by existing HAIC evaluation methodologies, while it aims at integrating and complement them.
+While the framework has not yet been applied in real-world scenarios, it is designed to be adaptable and to serve as a reference for future empirical validation.
+Section 2 presents a literature review on HAIC evaluation methods, summarizing different approaches, key metrics, and domain-specific challenges.
+This review highlights gaps in existing evaluation strategies, particularly their inability to fully capture the dynamic and reciprocal nature of HAIC interactions.
+Based on the finding of this review and the shortcoming identified we introduce an structured evaluation framework [bolman2013reframing] that systematically assesses HAIC across multiple domains (Sections 3– 4) .
+This framework integrates key evaluation factors such as goals, interaction quality, and task allocation, ensuring a standardized and adaptable assessment tool for practitioners and researchers.
+1.1 Background
+The rapid advancement of Artificial Intelligence (AI) has transformed multiple domains, leading to a shift from traditional Human-Machine Interaction (HMI) toward Human-AI Collaboration (HAIC) [1].
+Unlike conventional HMI, which emphasizes usability and task performance [2, 3], HAIC involves a symbiotic partnership where humans and AI work together to achieve shared goals.
+This paradigm shift has demonstrated immense potential across diverse sectors, including manufacturing [4, 5], financial services [6, 7], healthcare [8, 9], and education [10].
+However, despite its widespread adoption, evaluating HAIC remains an unresolved challenge.
+Traditional assessment frameworks, which focus on efficiency and accuracy, are insufficient to capture the complexity of human-AI interaction, adaptability, and trust.
+This limitation motivates the need for a comprehensive review of HAIC evaluation methodologies (Section 2) and the development of a structured framework that accounts for these multidimensional aspects (Sections 3– 4).
+1.2 Challenges
+While HAIC offers numerous opportunities, evaluating its effectiveness remains a complex and unresolved challenge.
+Recent work by Woelfle et al.
+(2024) [11] highlights the limitations of traditional evaluation metrics such as task performance, response time, and user satisfaction.
+Their study emphasizes the need for more comprehensive benchmarks that account for the dynamic and reciprocal nature of HAIC interactions, incorporating metrics like evidence appraisal quality, trust, and adaptability.
+Such advancements aim to better reflect the multifaceted and collaborative dimensions of human-AI partnerships.
+One of the primary challenges lies in assessing the individual and joint contributions of humans and AI systems to collaborative outcomes.
+Unlike traditional human-machine paradigms, HAIC involves a deeper integration of human intuition and AI capabilities.
+Evaluating this synergy requires novel metrics that effectively capture the mutual influences between collaborators.
+The unique challenges posed by creative and linguistic domains further highlight the limitations of traditional evaluation methods.
+Applications such as Large Language Models (LLMs) and generative AI in the arts demonstrate the transformative potential of AI to enhance creativity and human expression.
+However, evaluating these systems requires specialized metrics that move beyond task-based measures.
+For instance, assessing the interplay of human and AI contributions in artistic creation or the ethical implications of AI-generated content demands methodologies that capture qualitative and contextual nuances.
+While this study provides a foundational framework for HAIC evaluation, exploring the specific requirements of creative and linguistic domains is an essential avenue for future research.
+Additionally, measuring the quality of human-AI interactions remains a critical yet underexplored area.
+Interaction quality encompasses shared decision-making processes, communication dynamics, and mutual adaptability.
+Evaluating the ability of both the AI system and human participants to respond to evolving circumstances is crucial for understanding the resilience and effectiveness of HAIC systems in dynamic and high-stakes environments.
+1.3 Proposed Approach
+Drawing inspiration from established frameworks [12, 2], our approach emphasizes ethical considerations, adaptability, and domain-specific customization.
+The proposed framework provides a comprehensive structure for assessing HAIC across various modes of interaction, from AI-centric to symbiotic collaborations, integrating quantitative and qualitative measures.
+Furthermore, it aims to assist researchers and practitioners in identifying existing gaps and opportunities for improving collaboration between humans and AI.
+By incorporating domain-specific insights, the framework is designed to address practical challenges and applications across fields such as healthcare, finance, and education, thereby demonstrating its versatility and relevance in diverse contexts.
+1.4 Contributions
+This paper makes the following key contributions:
+  • A review of existing HAIC evaluation methodologies, identifying key gaps and challenges that limit their applicability across diverse domains.
+  • The introducion a structured evaluation framework for HAIC, integrating both quantitative and qualitative metrics to assess performance, interaction quality, and task allocation.
+  • The framework is domain-agnostic, designed to support AI-centric, human-centric, and symbiotic collaboration models, offering a flexible structure applicable to various application areas.
+2 Literature Review
+This section first examines the transition from HMI to HAIC, highlighting key evaluation challenges.
+It then explores the fundamental elements that define HAIC interactions, providing the foundation for developing a structured evaluation framework.
 2.1 From Human-Machine Interaction to Human-AI Collaboration
-Given the rapid pace of advancements in Human-AI Collaboration (HAIC), a comprehensive understanding of the current landscape is essential.
-While foundational research in Human-Machine Interaction (HMI) provides valuable insights into human-technology interactions, it often falls short when applied to HAIC evaluations.
-Traditional HMI methods, as exemplified in the works of Nielsen [20], Shneiderman [21], and Norman [22], primarily focus on assessing the usability and effectiveness of tools or interfaces designed to augment human performance.
-These methods often rely on metrics such as task completion time, error rates, and user satisfaction to gauge the success of human-machine interactions.
-However, these metrics are not always sufficient to capture the complexities of HAIC, where the relationship between humans and AI is more dynamic and collaborative.
-For example, traditional metrics may not adequately assess the level of trust a human places in an AI system’s recommendations or the effectiveness of shared decision-making processes where both the human and AI contribute to the final outcome [23].
-Additionally, these metrics may not fully capture the adaptability of AI systems, which is crucial for HAIC as the AI needs to learn and adjust its behavior based on human input and changing circumstances [24].
-To effectively evaluate HAIC, we must move beyond the limitations of traditional HMI approaches.
-We need to consider not just the performance of the individual components (human and AI) but also the quality of their collaboration, the fluidity of their interactions, and the achievement of shared goals.
-This necessitates a broader range of metrics and methodologies that can capture the nuances of these complex partnerships.
-2.2 Elements of HAIC
-Human-AI Collaboration (HAIC) is the cooperative partnership between individuals and artificial intelligence (AI) systems in order to accomplish common goals.
-Several key elements define the landscape of HAIC:
-  • Tasks: HAIC systems can be designed to tackle a wide array of tasks, from complex decision-making to creative endeavors and knowledge work.
-The nature of these tasks often dictates the level and type of collaboration required between humans and AI [25, 26].
-For example, in the medical field, an AI system can analyze medical images to identify potential anomalies, while a radiologist interprets the results and makes a final diagnosis.
-  • Goals: HAIC is driven by shared goals, which can be both individual and collective.
-Individual goals might encompass AI objectives like improving efficiency and accuracy or human aims like skill enhancement and knowledge acquisition.
-Collective goals focus on achieving overarching objectives that benefit from the combined capabilities of both humans and AI [27].
-For example, in a healthcare setting, an AI system could be designed to examine medical images and identify initial indications of illnesses like cancer (individual goal).
-The human radiologist review’s the the AI’s findings considers patient history, and makes the final diagnosis (individual goal).
-The collective goal in this scenario is to improve diagnostic accuracy and ensure timely and effective patient care, thereby enhancing overall health outcomes.
-  • Interaction: The success of HAIC heavily relies on effective communication and feedback mechanisms between humans and AI.
-The quality of this interaction determines how well each party understands the other’s intentions, capabilities, and limitations [17, 28].
-For instance, an AI chatbot for customer service needs to understand and respond to customer queries effectively, adapting its communication style based on user feedback.
-  • Task Allocation: Assigning tasks appropriately based on the strengths and weaknesses of both humans and AI is critical.
-Dynamic task allocation, where responsibilities shift based on real-time needs, is often a hallmark of successful HAIC [3].
-A good example is in autonomous vehicles, where AI systems may be responsible for navigation under normal conditions, but the human driver takes control when faced with unexpected situations.
+The rapid advancement of HAIC necessitates a broader evaluation perspective that goes beyond traditional HMI usability assessments [13, 14, 15].
+HMI methods have primarily been designed to measure user experience and task performance, relying on metrics such as response time, accuracy, and efficiency.
+However, HAIC introduces new challenges that extend beyond interface usability, requiring metrics that capture interaction quality, collaboration effectiveness, and trust.
+Okamura and Yamada (2020) [16] emphasize the importance of system transparency in maintaining trust calibration between humans and AI agents.
+Over-trusting AI can lead to automation bias, while insufficient trust may result in underutilization.
+Similarly, Yue et al.
+(2023) [17] examine how different human-AI collaboration models influence user perceptions and adoption, reinforcing the need for evaluation methods that incorporate trust, interpretability, and adaptability.
+2.2 Core Elements and Modes of HAIC
+Human-AI Collaboration (HAIC) is defined by the cooperative partnership between individuals and AI systems to accomplish shared objectives.
+Understanding its key elements is essential for designing effective evaluation methodologies.
+  • Tasks: HAIC systems support diverse tasks, ranging from automated decision-making to creative problem-solving.
+The nature of a given task dictates the required level of collaboration between humans and AI [18, 19].
+For instance, in healthcare, an AI system may analyze medical images, while a radiologist interprets and integrates results into a final diagnosis.
+  • Goals: Collaboration in HAIC is driven by individual and collective goals.
+AI objectives often include enhancing efficiency and accuracy, while human goals may involve skill augmentation and contextual reasoning [20].
+Continuing with the radiology example, the AI assists in accurate disease detection, while radiologists focus on clinical decision-making and patient-centered care.
+  • Interaction: The effectiveness of HAIC is determined by communication quality and feedback mechanisms [21, 22].
+Misalignment in expectations or miscommunication can lead to ineffective collaboration.
+For instance, AI-driven customer support chatbots must interpret user queries and refine responses dynamically to ensure effective problem resolution.
+  • Task Allocation: Dynamic task allocation is essential in HAIC, ensuring that responsibilities shift based on real-time conditions and expertise distribution [23].
+For example, in advanced driver assistance systems, AI controls navigation under normal conditions, but the human driver intervenes when unexpected obstacles arise.
 Figure 1 visually represents the key elements of HAIC, which are important for developing effective evaluation frameworks.
@@
 Figure 1: Core elements of Human-AI Collaboration (HAIC) and examples of their application.
-2.3 Modes of Human-AI Collaboration
 One can use any of the HAIC elements, i.e., tasks, goals, types of interaction, and task allocation to differentiate between different HAIC methods.
@@
 In the Human-centric mode, humans retain the primary decision-making authority, utilizing AI as an augmentative tool to enhance human capabilities without superseding the human role.
-This mode values human intuition and oversight, with AI employed to manage data-intensive or repetitive tasks.
+This mode values human intuition and oversight, with artificial intelligence employed to manage repetitive or data-intensive tasks.
 It optimizes the unique cognitive abilities of humans for strategic and ethical decision-making while leveraging AI to augment human capabilities.
@@
 [30] further illustrate how symbiotic HAIC can revolutionize creative fields and managerial roles by combining human creativity with AI’s analytical strengths to foster innovation and improve decision-making.
-A key example of the symbiotic mode is the Learning to Defer paradigm.
-This paradigm designs AI systems to understand their own limitations, knowledge, and expertise.
-Whether confronted with ambiguity or circumstances in which their confidence in a decision is low, the AI system proactively defers to human judgment  [36].
-Several studies have explored this concept in depth.
-Mozannar and Sontag (2020) [37] proposed a Bayesian framework that emphasizes the integration of human and machine predictions.
-It is crucial to consider potential biases in both agents.
-This approach aligns with a collaborative decision-making approach because it strives for optimization.
-It is possible to make decisions by leveraging the strengths of both humans and AI.
-In a similar way, Mao et al.
-(2023) [38] extended this concept to scenarios with multiple experts, highlighting the flexibility required in real-world HAIC systems, where multiple sources of expertise may be available.
+A key example of the symbiotic mode is the Learning to Defer (L2D) paradigm, where AI systems are designed to recognize their own limitations and defer decisions to human judgment when necessary.
+This approach enhances collaborative decision-making by leveraging the strengths of both humans and AI.
+Recent studies have expanded on this concept.
+Alves et al.
+(2024) [31] introduced a cost-sensitive L2D framework that considers scenarios with multiple experts and workload constraints, emphasizing the need for flexibility in real-world Human-AI Collaboration (HAIC) systems.
+Additionally, Hemmer et al.
+(2022) [32] proposed an approach that trains a classification model to complement the capabilities of multiple human experts.
 Their work underscores the adaptability of the HAIC systems’ participants, showcasing the ability to dynamically modify task allocation according to the availability and expertise of various agents.
@@
 In this section, we looked at the fundamentals of Human-AI Collaboration, such as its main components and the various modes of operation.
-Evaluating HAIC is critical to ensuring that these systems are successful, efficient, and ethical.
+The evaluation of HAIC is critical to ensure that these systems are successful, efficient, and ethical.
 The following part looks into the literature on HAIC evaluation, reviewing existing approaches and identifying gaps that suggest areas for further research.
-3 Evaluation of Human-AI Collaboration in the Research Literature
+2.3 Evaluation Approaches in HAIC
 The evaluation of Human-AI Collaboration (HAIC) is essential to understanding the efficacy of these systems, identifying areas for improvement, and ultimately unlocking their full potential.
@@
   • To effectively address the intricate and diverse aspects of HAIC in different sectors.
-{adjustwidth}
--1cm-1cm Reference Domain AI Technology/Method Qn Ql Mix Additional Info Vollmuth et al.
-(2023) [41] Healthcare (Oncology) Image Analysis Accuracy is measured by sensitivity, precision, recall, F1-score.
-Focuses on human specialists to evaluate AI and HAIC performance.
-Van Leeuwen et al.
-(2022) [42] Healthcare (Radiology) Image Analysis Emphasizes efficiency using sensitivity, specificity, and outcome metrics.
-Timmons et al.
-(2023) [43] Healthcare (Mental Health) Not specified Investigates bias through interviews and case studies.
-Farić et al.
-(2024) [44] Healthcare (Radiology) Diagnostic Support Systems Assesses early AI use via qualitative feedback from radiologists.
-Calisto et al.
-(2022) [45] Healthcare (Breast Cancer) Medical Image Analysis Focuses on human-AI interaction through qualitative methods.
-Rezwana et al.
-(2023) [46] Design Generative AI Develops interaction design framework using user feedback.
-Tschandl et al.
-(2020) [10] Healthcare (Dermatology) Image Analysis Uses quantitative success metrics in high-stakes environments.
-Siu et al.
-(2021) [47] Game-Based Reinforcement Learning Uses game performance metrics in Hanabi game as a testbed.
-Reddy et al.
-(2021) [48] Healthcare Not specified Proposes a qualitative framework focusing on ethics and user perspectives.
-Sharma et al.
-(2023) [49] Mental Health Reinforcement Learning Measures conversational empathy improvements through qualitative feedback.
-Yik et al.
-(2024) [50] General (Benchmarking) Neuromorphic Computing Provides quantitative framework for unique AI context.
-Rastogi et al.
-(2023) [51] General (Auditing) LLM Examines auditing collaboration using qualitative assessments.
-Zhao et al.
-(2023) [52] Multimodal Interaction LLM with Visual Input Studies interaction through multimodal input performance metrics.
-Makarius et al.
-(2020) [1] General (Business) Not specified Emphasizes social factors using qualitative methods.
-Aleid et al.
-(2023) [53] Healthcare (Radiology) Image Analysis Focuses on accuracy using quantitative measures.
-Al-Fatlawi et al.
-(2023) [54] Finance (Fraud) Genetic Algorithms Evaluates detection rates (true/false positives).
-Yang et al.
-(2022) [55] General (Teaming) Behavior Modeling Uses quantitative and qualitative data to model human behavior for better interaction.
-Aher et al.
-(2023) [56] Simulation-Based Large Language Models Simulates human behavior using quantitative metrics.
+Table 1: Examples of Evaluation of Human-AI Collaboration Across Sectors
+Reference Domain AI Technology/Method Qn Ql Mix Additional Info
+Vollmuth et al. (2023) [35] Healthcare (Oncology) Image Analysis Accuracy is measured by sensitivity, precision, recall, F1-score. Focuses on human specialists to evaluate AI and HAIC performance.
+Van Leeuwen et al. (2022) [36] Healthcare (Radiology) Image Analysis Emphasizes efficiency using sensitivity, specificity, and outcome metrics.
+Timmons et al. (2023) [37] Healthcare (Mental Health) Not specified Investigates bias through interviews and case studies.
+Farić et al. (2024) [38] Healthcare (Radiology) Diagnostic Support Systems Assesses early AI use via qualitative feedback from radiologists.
+Calisto et al. (2022) [39] Healthcare (Breast Cancer) Medical Image Analysis Focuses on human-AI interaction through qualitative methods.
+Rezwana et al. (2023) [40] Design Generative AI Develops interaction design framework using user feedback.
+Tschandl et al. (2020) [8] Healthcare (Dermatology) Image Analysis Uses quantitative success metrics in high-stakes environments.
+Siu et al. (2021) [41] Game-Based Reinforcement Learning Uses game performance metrics in Hanabi game as a testbed.
+Reddy et al. (2021) [42] Healthcare Not specified Proposes a qualitative framework focusing on ethics and user perspectives.
+Sharma et al. (2023) [43] Mental Health Reinforcement Learning Measures conversational empathy improvements through qualitative feedback.
+Yik et al. (2024) [44] General (Benchmarking) Neuromorphic Computing Provides quantitative framework for unique AI context.
+Rastogi et al. (2023) [45] General (Auditing) LLM Examines auditing collaboration using qualitative assessments.
+Zhao et al. (2023) [46] Multimodal Interaction LLM with Visual Input Studies interaction through multimodal input performance metrics.
+Makarius et al. (2020) [1] General (Business) Not specified Emphasizes social factors using qualitative methods.
+Aleid et al. (2023) [47] Healthcare (Radiology) Image Analysis Focuses on accuracy using quantitative measures.
+Al-Fatlawi et al. (2023) [48] Finance (Fraud) Genetic Algorithms Evaluates detection rates (true/false positives).
+Yang et al. (2022) [49] General (Teaming) Behavior Modeling Uses quantitative and qualitative data to model human behavior for better interaction.
+Aher et al. (2023) [50] Simulation-Based Large Language Models Simulates human behavior using quantitative metrics.
 Verheijden & Funk [51] (2023) Design Generative AI Enhances design processes through qualitative co-creation feedback.
-Hauptman et al.
-(2023) [58] General (Teaming) Adaptive AI Investigates adaptability perceptions through qualitative assessments.
-Sankaran et al.
-(2022) [59] Manufacturing Recurrent Neural Networks Measures process performance using quantitative metrics.
+Hauptman et al. (2023) [52] General (Teaming) Adaptive AI Investigates adaptability perceptions through qualitative assessments.
+Sankaran et al. (2022) [53] Manufacturing Recurrent Neural Networks Measures process performance using quantitative metrics.
 Massaro (2022) [54] Manufacturing (Safety) Artificial Neural Networks Evaluates safety and risk through quantitative measures.
 Oravec (2023) [55] Education Large Language Models Discusses academic integrity using qualitative methods.
-El-Zanfaly et al.
-(2022) [62] Design/Art Physical Interface Focuses on creativity and interaction through qualitative feedback.
+El-Zanfaly et al. (2022) [56] Design/Art Physical Interface Focuses on creativity and interaction through qualitative feedback.
 Arias-Rosales (2022) [57] Design/Art Shape Generation Combines quantitative shape metrics with qualitative value assessments.
-Dziorny et al.
-(2022) [64] Healthcare (ICU) Decision Support Systems Examines high-stakes challenges using qualitative feedback.
-Vössing et al.
-(2022) [65] General (Forecasting) LightGBM, Neural Networks Investigates AI transparency impact using mixed trust/performance metrics.
-Chakravorti et al.
-(2022) [66] Finance Prediction Models Combines detection rates with qualitative trust assessments for AI markets.
+Dziorny et al. (2022) [58] Healthcare (ICU) Decision Support Systems Examines high-stakes challenges using qualitative feedback.
+Vössing et al. (2022) [59] General (Forecasting) LightGBM, Neural Networks Investigates AI transparency impact using mixed trust/performance metrics.
+Chakravorti et al. (2022) [60] Finance Prediction Models Combines detection rates with qualitative trust assessments for AI markets.
 Huang & Rust (2022) [61] Marketing Machine Learning Develops collaborative AI framework using mixed methods.
-Sachan et al.
-(2024) [68] Finance Evidential Reasoning Reduces decision noise through quantitative measures.
-Basu et al.
-(2021) [69] Finance (ICO) Linear Regression Studies decision-making using mixed quantitative and qualitative data.
+Sachan et al. (2024) [62] Finance Evidential Reasoning Reduces decision noise through quantitative measures.
+Basu et al. (2021) [63] Finance (ICO) Linear Regression Studies decision-making using mixed quantitative and qualitative data.
 Dikmen & Burns (2022) [64] Finance (Peer-to-Peer Lending) Explainable AI Examines trust and performance using mixed domain knowledge assessments.
-Ma et al.
-(2024) [71] General (Decision-Making) Large Language Models Proposes deliberative AI using mixed decision-making metrics.
+Ma et al. (2024) [65] General (Decision-Making) Large Language Models Proposes deliberative AI using mixed decision-making metrics.
 Alon-Barkat & Busuioc (2023) [66] Public Sector Regression Models Explores public sector interactions through mixed methods.
 Nikolopoulou (2024) [67] Education Generative AI Assesses generative AI’s impact on pedagogy using mixed methods.
-Table 1: Generalized Evaluation of Human-AI Collaboration Across Sectors
-3.1 Advancements in Evaluation Frameworks
-Human-AI Collaboration is being judged by a more traditional metrics. These include ethical concerns, the ability to scale and the ability to adapt different domains.
-Dellermann et al.
-(2021) [25] propose a classification system for hybrid intelligence, emphasizing the importance of understanding the nature of tasks and the dynamic relationship between humans and AI.
-To categorize different types of HAIC systems, their framework takes into account factors such as level of autonomy, task interdependence, and degree of human control.
-This classification scheme serves as a foundation for developing tailored evaluation approaches that address the unique characteristics of each type of collaboration.
+2.3.1 Advancements in Tools, Methodologies, and Real-Time Collaboration for HAIC
+Human-AI Collaboration (HAIC) is increasingly evaluated using both traditional and innovative metrics, with an emphasis on ethical concerns, scalability, and adaptability across domains.
+Recent advancements in tools and methodologies are reshaping the evaluation landscape, offering novel ways to understand and optimize human-AI interactions.
+Fischer (2023) [68] envisions the future of human-AI collaboration as a deeply intertwined partnership, where AI systems are designed not only to assist with tasks but also to foster positive emotional and social interactions.
+His work explores scenarios where AI colleagues are capable of understanding and responding to human emotions, aiming to enhance collaboration by creating a sense of trust and mutual satisfaction.
+This forward-looking perspective emphasizes the need for evaluation frameworks that address not only technical performance but also the social and emotional dimensions of human-AI interactions.
+Fischer’s work highlights the importance of incorporating human-centered design principles into AI systems to promote effective collaboration and user well-being.
 The incorporation of ethical factors into evaluation frameworks is a growing trend, as evidenced by the work of Huang & Rust (2022) [61] in marketing and Nasir, Khan, & Bai (2023) [69] in healthcare.
@@
 This framework outlines a step-by-step process for defining evaluation objectives, selecting appropriate metrics (e.g., accuracy, efficiency, user satisfaction), designing experimental protocols, and analyzing data to draw valid conclusions about the effectiveness of HAIC.
-3.2 Advancements in Evaluation Tools and Methodologies: Enabling Real-time Insights and Adaptive Collaboration
-Technological advancements are transforming the landscape of HAIC evaluation, enabling real-time monitoring, adaptive collaboration, and deeper insights into human-AI interactions.
-AI-powered analytics and real-time monitoring tools, discussed by Wang et al.
-(2019) [77], are enhancing the precision and responsiveness of HAIC evaluations.
-These tools can automatically track key performance indicators (KPIs) such as completion time, error rates, and user interactions, providing immediate feedback on collaboration efficacy and enabling timely interventions for improvement.
-Advancements in interactive machine learning (iML) interfaces, explored by Saha et al.
-(2023) [78], are fostering grater transparency and user understanding.
-By providing explanations for AI-generated recommendations and allowing users to adjust system parameters, iML interfaces can improve user comprehension, trust, and satisfaction, ultimately contributing to a more effective and collaborative partnership between humans and AI.
+Technological advancements further enable real-time monitoring and adaptive collaboration in HAIC.
+Recent studies, such as Woelfle et al.
+(2024) [11], emphasize the importance of benchmarking tools that evaluate human-AI collaboration through diverse performance metrics.
+These metrics extend beyond traditional KPIs like task completion time and error rates, incorporating dimensions such as evidence appraisal quality and the adaptability of AI systems to human inputs.
+Such tools provide actionable feedback on collaboration efficacy, fostering dynamic adaptability and enhancing the overall effectiveness of HAIC systems.
+Interactive machine learning (iML) interfaces, explored by Saha et al.
+(2023) [72], represent a significant leap in fostering transparency and user engagement.
+By providing explanations for AI-generated recommendations and allowing users to adjust system parameters, iML interfaces enhance user comprehension, trust, and satisfaction.
 Saha et al.
-[78] evaluate the effectiveness of iML interfaces using a combination of user surveys, interviews, and task-based performance metrics.
-3.3 Analysis of Existing Approaches
+evaluate these interfaces using a combination of user surveys, interviews, and task-based performance metrics, demonstrating their potential to facilitate effective human-AI partnerships.
+Together, these advancements underscore the importance of integrating ethical considerations, real-time analytics, and interactive tools to develop comprehensive evaluation frameworks for HAIC.
+These tools and methodologies not only enhance our understanding of human-AI collaboration but also pave the way for more adaptive, transparent, and ethical partnerships between humans and AI systems.
+2.3.2 Analysis of Existing Approaches
 In Table 1, we show a summary of Human-AI Collaboration (HAIC) evaluation practices we examined for this analysis.
@@
 Reddy et al.
-(2021) [48] suggest a qualitative framework for evaluating the use of AI in healthcare, with a focus on ethical issues and user perspectives.
+(2021) [42] suggest a qualitative framework to evaluate the use of AI in healthcare, with a focus on ethical issues and user perspectives.
 Their framework includes structured interviews with healthcare providers and patients, and thematic analysis to identify key concerns such as data privacy, trust in AI diagnoses, and the perceived impact on the patient-provider relationship.
-Their study underscores the necessity for AI systems to align with ethical standards and user expectations to be effectively integrated into healthcare practices.
+Their study underscores the need for AI systems to align with ethical standards and user expectations to be effectively integrated into healthcare practices.
 By capturing the nuanced perspectives of users and stakeholders, these qualitative studies offer a rich understanding of the human factors that shape the adoption and impact of HAIC.
-However, the subjective nature of qualitative data can pose challenges in terms of generalizability and require substantial resources for data collection and analysis.
-The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
-The limitations of purely quantitative or qualitative approaches have spurred the adoption of mixed-methods evaluations, aiming to capture both objective performance and subjective experiences in a comprehensive manner.
+However, the subjective nature of qualitative data can pose challenges in terms of generalizability and requires substantial resources for data collection and analysis.
+2.3.3 The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
+The limitations of relying solely on quantitative or qualitative approaches have led to the adoption of mixed-methods evaluations, which integrate objective performance metrics with subjective experiences.
+This approach provides the flexibility needed to tailor evaluations to specific domains while balancing the rigor of quantitative data with the depth of qualitative insights.
 Domain-Specific Applications
-In the realm of design and art, Arias-Rosales (2022) [63] combines quantitative shape generation metrics with qualitative assessments of perceived value by designers.
-Their study uses quantitative metrics to measure the efficiency and novelty of AI-generated shapes, while also conducting interviews and surveys with designers to gather insights on the aesthetic value and usability of these shapes.
-This approach revealed that while AI-generated designs were technically innovative, the perceived value varied significantly among designers, highlighting the importance of subjective aesthetic judgment in evaluating creative outputs.
+In creative domains such as design and art, Arias-Rosales (2022) [57] combined quantitative shape generation metrics with qualitative assessments of designers’ perceived value.
+This dual approach highlighted that while AI-generated designs were technically innovative, their subjective aesthetic value varied significantly among designers, underlining the critical role of qualitative judgment in evaluating creative outputs.
 Similarly, in finance, Chakravorti et al.
-(2022) [66] integrate quantitative detection rates with qualitative assessments of trust and perception to evaluate the efficacy of AI prediction markets.
-They employ quantitative methods to analyze the accuracy and reliability of AI predictions in financial markets, alongside qualitative interviews and surveys with traders and financial analysts.
-The study found that despite high detection rates, the trust in AI predictions varied, with qualitative data uncovering concerns about the transparency and interpretability of AI algorithms.
+(2022) [60] evaluated AI prediction markets by pairing detection rates with trader and analyst perceptions.
+Their mixed-methods approach revealed that high prediction accuracy did not necessarily translate into trust, exposing concerns about transparency and interpretability.
 Generalizable Frameworks
-Yang et al.
-(2022) [55] look into how AI can model human behavior using both quantitative and qualitative data.
-They use quantitative metrics such as task completion rates and error frequencies to evaluate AI performance, while also gathering qualitative data through user interviews and behavioral observations.
-Vössing et al.
-(2022) [79] study includes quantitative measures of system accuracy and user performance metrics, coupled with qualitative trust assessments obtained from user interviews and surveys.
-The combination of these methods revealed that transparency significantly impacts user trust and acceptance, even when the AI system’s performance is objectively high.
-Huang & Rust (2022) [67] propose a mixed-methods framework for collaborative AI in marketing, acknowledging the importance of both performance and ethical considerations.These mixed-methods techniques provide the possibility for a comprehensive and contextual evaluation of HAIC, recognizing the intricate relationship between objective measurements and subjective perceptions.
-They employ quantitative analytics to measure marketing campaign performance metrics, such as click-through rates and conversion rates, while also using qualitative methods like focus groups and in-depth interviews to assess ethical concerns and user satisfaction.
-This dual approach provides a comprehensive evaluation of the AI’s effectiveness and its alignment with ethical standards.
-3.4 The domain-specific challenges necessitate a tailored evaluation
-The application domains for HAIC are diverse, each presenting unique challenges that necessitate tailored evaluation approaches.
-These challenges stem not only from the technical complexities of integrating AI into different workflows, but also from the varying priorities and expectations of stakeholders in different sectors.
-As a result, the evaluation methods employed across HAIC research vary significantly, reflecting the need to adapt assessment strategies to the specific context of each domain.
+Mixed methods also demonstrate utility across domains.
+For instance, Yang et al.
+(2022) [49] modeled human behavior by combining task completion rates and error frequencies with user interviews, while Vössing et al.
+(2022) [73] integrated system accuracy metrics with trust assessments, uncovering the significant impact of transparency on user acceptance.
+In marketing, Huang & Rust (2022) [61] employed a mixed-methods framework that combined performance metrics, such as click-through rates, with ethical and user satisfaction insights.
+Together, these studies illustrate how mixed methods enable comprehensive evaluations that capture both technical performance and contextual user perceptions.
+2.3.4 Tailored Evaluation for Domain-Specific Challenges
+HAIC’s diverse application domains present unique challenges that demand tailored evaluation approaches, reflecting both technical complexities and the distinct priorities of stakeholders.
+As discussed in the Appendix A of this work, the evaluation of creative and linguistic AI requires tailored methodologies due to its inherent complexities.
 Healthcare: Prioritizing Accuracy, Safety, and Ethical Considerations
-In the healthcare sector, ensuring diagnostic accuracy and patient safety is paramount.
-To find out how well AI-assisted tools work in oncology and radiology, studies like Vollmuth et al.
+In the healthcare, diagnostic accuracy and patient safety are critical.
+Studies like Vollmuth et al.
 (2023) [35] and Van Leeuwen et al.
-(2022) [42] use numbers like sensitivity, specificity, and outcome improvements.
-However, the high-stakes nature of medical decisions and the potential for unintended consequences necessitate additional layers of evaluation.
-Qualitative approaches like interviews and focus groups (Farić et al., 2024 [14]; Calisto et al., 2022 [45]) can capture user perceptions, identify potential biases (Timmons et al., 2023 [43]), and ensure that AI systems are implemented in a way that aligns with ethical principles and patient values (Reddy et al., 2021 [48]).
+(2022) [36] use metrics like sensitivity, specificity but the high-stakes nature of medical decisions require additional qualitative approaches like interviews and focus groups (Farić et al., 2024 [74]; Calisto et al., 2022 [39]) to capture user perceptions, identify biases (Timmons et al., 2023 [37]), and ensure alignment with ethical principles (Reddy et al., 2021 [42]).
 Creative Arts: Balancing Creativity, User Experience, and Ethical Impact
-In the creative arts, HAIC evaluation often requires a different approach.
-While quantitative metrics can measure the technical performance of AI systems, such as the speed or accuracy of generating artistic outputs, qualitative methods such as interviews, user feedback, and case studies are crucial for comprehending the influence of AI on creativity, collaboration, and the overall design process (Rezwana et al., 2023 [46]; Verheijden & Funk, 2023 [57]).
-Additionally, ethical considerations, such as the potential for AI to perpetuate biases or undermine the role of human creativity, necessitate careful evaluation through qualitative assessments and critical analysis.
-Moreover, the cooperative aspect of HAIC plays a crucial role in enhancing human creativity, since AI’s data-driven insights stimulate inventive solutions [80].
+In creative arts, quantitative metrics (e.g.
+speed, output quality) are complemented byqualitative methods such as interviews and case studies (Rezwana et al., 2023 [40]; Verheijden & Funk, 2023 [51]) to asses AI’s influence on creativity and collaboration.
+Ethical concerns regarding bias and the potential undermining of human creativity, underscore the need for a balanced evaluation [75].
 Finance: Mitigating Risks, Ensuring Trust, and Balancing Quantitative and Qualitative Insights
-The financial sector employs a mix of quantitative and qualitative methods to address HAIC’s unique challenges in this domain.
-Quantitative metrics like detection rates and false positives are crucial for evaluating the effectiveness of AI in fraud detection (Al-Fatlawi et al., 2023 [54]).
-However, understanding the impact of AI on decision-making processes, investor trust, and market stability requires qualitative insights through surveys, interviews, and case studies (Chakravorti et al., 2022 [66]).
-Mixed-methods approaches, which combine quantitative and qualitative data, offer a more comprehensive understanding of the complex interactions between humans, AI, and financial markets.
-In summary, the review of evaluation approaches demonstrates significant variation in collaboration methods, domain-specific challenges, and methodological approaches.
-However, one significant obstacle remains: the lack of a comprehensive, mixed-methods evaluation framework that covers a wide range of industries.
-This gap restricts our ability to completely evaluate HAIC’s effectiveness and maximize its potential across domains.
-The Path Forward: A Unified Evaluation Framework to Bridge the Gap in HAIC Assessment
-The analysis of existing approaches to HAIC evaluation reveals a landscape marked by diversity and fragmentation.
-Quantitative, qualitative, and mixed-methods approaches all give useful information, but the lack of a single, flexible framework makes it harder to get a full picture of how well HAIC works and what it can do in different fields.
-This gap is particularly evident when considering the multifaceted nature of HAIC.
-Quantitative metrics, while essential for measuring objective performance, often fail to capture the nuances of human-AI interaction dynamics, such as trust, satisfaction, and ethical considerations.
-Qualitative methods, on the other hand, can provide rich insights into these subjective experiences but may lack the rigor and generalizability of quantitative data.
-Therefore, it is imperative to develop a comprehensive, mixed-methods evaluation framework.
-This framework should be adaptable across a broad spectrum of industries, accommodating the unique characteristics and challenges of each while providing a cohesive approach to assessing the effectiveness, challenges, and opportunities of HAIC.
-By integrating objective performance metrics with subjective experiences, such a framework can offer a more nuanced understanding of how HAIC systems perform in real-world contexts and how they impact various stakeholders.
-This unified framework would not only address the current methodological fragmentation but also facilitate cross-study comparison, enhance the generalizability of findings, and promote standardized evaluation practices.
-It would enable researchers, practitioners, and policymakers to make informed decisions based on a more holistic understanding of the complex interactions between humans and AI.
-In healthcare, this framework could integrate measures of diagnostic accuracy with assessments of user trust and ethical considerations, ensuring that AI-assisted tools are both effective and aligned with patient values.
-In creative fields, it could incorporate quantitative metrics of creative output alongside qualitative assessments of the impact on artistic processes and collaboration.
-In finance, it could combine quantitative measures of risk and return with qualitative assessments of trust, transparency, and the impact on decision-making.
-By addressing the limitations of existing approaches and providing a standardized yet adaptable framework, we can unlock the full potential of HAIC across diverse domains.
-This would enable us to harness the power of AI in ways that augment human capabilities, foster innovation, and ultimately contribute to societal well-being.
-The development of such a framework is not merely an academic pursuit but a critical step towards ensuring the responsible and impactful integration of AI into human workflows.
+In finance quantitative metrics such as detection rates and false positives are crucial for fraud detection (Al-Fatlawi et al., 2023 [48]).
+However, understanding AI’s impact on decision-making and market stability also requires qualitative insights through surveys and interviews (Chakravorti et al., 2022 [60]).
+Mixed-methods approaches, provide a comprehensive view of human-AI interaction in financial markets.
+For a detailed discussion on the challenges specific to creative and linguistic AI, please refer to Appendix A.
+In summary, while evaluation approaches vary significantly across domains, a comprehensive, mixed-methods framework is still needed to sully assess HAIC’s effectiveness and potential.
+2.4 Towards a Unified Evaluation Framework
+While existing HAIC evaluation approaches offer valuable insights, they remain fragmented.
+A unified framework is needed to integrate objective performance metrics with subjective experiences across diverse domains.
+Such a framework would support cross-study comparison, enhance generalizability, and standardize evaluation practices.
+This framework should adapt to different industries, ensuring AI-assisted tools are both effective and aligned with human needs.
+In healthcare, it could integrate diagnostic accuracy with user trust; in creative fields, it could assess AI’s impact on artistic processes; and in finance, it could combine risk metrics with transparency evaluations.
+By addressing methodological gaps, this approach will help harness AI’s potential while ensuring its responsible integration into human workflows.
 3 A Structured Framework for Human-AI Collaboration
-4.1 Factors of Identifying the Human-AI Collaboration
-Based on the previous discussion (Section 2) the evaluation of HAIC is centered around three primary factors: Goals, Interaction, and Task Allocation.
-These factors are critical for understanding and enhancing HAIC systems.
-While tasks themselves are fundamental to HAIC, they are considered the context within which collaboration occurs, rather than a core factor of the evaluation framework.
-The nature of the task will influence how the other three factors are approached, but the tasks themselves are not directly evaluated by this framework.The literature highlights the importance of these factors in evaluating HAIC.
-Research by Li et al.
-[81] and Vossing et al.
-[65] underscores the significance of evaluating goals through metrics such as the learning curve and prediction accuracy.
-Studies by Hinsen et al.
-[82] and Amershi et al.
-[83] emphasize the need for user friendly communication methods and clear interaction mechanisms for successful HAIC.
-Additionally, Xiong et al.
-[84] stress the importance of measuring query efficiency and expertise utilization for effective task allocation.
-Following the identification of these primary factors, we further explore the subfactors and metrics that can be used to evaluate each factor.
-These subfactors and metrics are designed to provide a comprehensive understanding of the effectiveness, efficiency, and usability of HAIC systems.
-The structured framework presented here serves as a guide for researchers, practitioners, and policymakers to evaluate and optimize HAIC systems across various domains and applications.
+Unlike previous studies that focus on isolated aspects of HAIC evaluation, the proposed framework provides a structured approach by setting measurable goals, clarifying tasks and reporting metrics [bolman2013reframing], that integrates multiple evaluation dimensions, organized in three primary factors and specialized subfactors.
+The primary factors are:
+  • Goals – Evaluates individual and collective objectives within HAIC systems, with key metrics such as learning curves and prediction accuracy [76, 59].
+  • Interaction – Assesses communication mechanisms, feedback processes, adaptability, and trust. Effective communication and transparency are critical for ensuring human-AI trust and usability [77].
+  • Task Allocation – Examines role distribution, efficiency, and decision-making processes, ensuring that AI contributions complement human expertise in dynamic environments [78].
+Each factor is further divided into measurable subfactors and associated metrics, allowing for a robust quantitative and qualitative assessment of HAIC effectiveness.
+To enhance practical applicability, we develop a decision-tree-based evaluation approach.
+This structured method assists practitioners in selecting relevant evaluation criteria based on the nature of the HAIC system.
+The decision tree ensures flexibility, allowing for tailored assessments in various domains, including healthcare, finance, manufacturing, and education.
+By offering a holistic assessment tool, this framework addresses the key gaps identified in the literature and establishes a standardized methodology for evaluating HAIC in both research and industry applications.
 3.1 Subfactors & Metrics of the Framework
-To promote further analysis and evaluation of the different types of interactions in HAIC, the three primary factors are further analyzed into measurable subfactors, which can be evaluated using a set of metrics.
-These are presented in Table 2.
-Following we provide a detailed description of the subfactors and metrics associated with each primary factor.
-Please note that Table 3 summarizes all metrics discussed in this section, including their formal definition.
-{adjustwidth}
--1cm-1cm Primary Factor Subfactor Human-AI Collaboration Metrics for Evaluation Goals Individual Goals AI’s Goal: Learn Efficiently, Increase Accuracy.
+Table 2 summarizes the detailed breakdown of factors to subfactors and their associated metrics.
+Table 2: Factors and Subfactors of Human-AI Collaboration with Associated Metrics for Evaluation
+Primary Factor Subfactor Human-AI Collaboration Metrics for Evaluation
+Goals Individual Goals AI’s Goal: Learn Efficiently, Increase Accuracy.
 Human’s Goal: Provide Effective Teaching, Achieve Task Objectives.
@@
 Interaction Communication Methods Query and Response Mechanism.
-* Ease of Use (user satisfaction with communication methods), * Clarity of Communication (understandability of queries and responses) [82, 83] Feedback Mechanisms Human feedback on AI queries, AI adjustments based on feedback.
-* Frequency of Feedback (number of feedback instances), * Feedback Quality (relevance and specificity of feedback), * Feedback Impact (measured change in AI performance due to feedback) [83, 88, 82] Adaptability AI’s ability to adapt query strategy and learning based on human input.
-* Strategy Adaptation Effectiveness (improvement in query selection based on feedback), * Adaptability Score (composite metric of flexibility and responsiveness) [82, 89, 90] Trust and Safety Trust in AI system, ensuring user safety.
-* Trust Score (level of user trust in AI system) [91], * Safety Incidents (number of safety-related issues) [92], * Error Reduction Rate (decrease in errors due to AI assistance) [4], * Confidence (level of user trust in AI recommendations) [93] Task Allocation Complementarity AI handles data analysis and uncertainty estimation; Humans provide expertise and clarification.
-* Query Efficiency (amount of information gained per query), * Expertise Utilization (degree to which human expertise is leveraged) [84] Flexibility Adjusting query strategy based on interaction dynamics.
-* Adaptability Score (same as in Interaction), * Dynamic Task Allocation (ability to adjust task distribution based on real-time needs) [3] Efficiency Streamlining the learning process through targeted queries.
-* Model Improvement Rate (rate of improvement in AI performance) [65, 84], * Learning Curve (same as in Goals), * Resource Utilization (efficiency in using computational resources) [65] Responsiveness AI’s ability to rapidly adjust to new inputs and feedback.
-* Response Time (time taken to respond to queries or feedback), * Real-time Performance (ability to handle time-sensitive tasks) [82, 94] Collaborative Decision Making Joint decisions on data labeling and correction of AI’s model.
-* Impact of Corrections (measured improvement in system performance through human intervention) [95, 96], * Decision Effectiveness (impact of decisions on overall task success) [97, 98] Continuous Learning AI’s ability to learn from human input and adjust its model.
-* Model Improvement Rate (same as in Efficiency), * Learning Curve (same as in Goals), * Knowledge Retention (ability to retain learned information over time) [99] Mutual Support AI’s ability to provide support to humans in the learning process; Humans support by providing information.
-* Error Reduction Rate (decrease in errors due to AI assistance) [4], * Confidence (level of user trust in AI recommendations) [100], * Task Completion Time (with and without AI support) [101] Robustness Ensuring AI system’s robustness and reliability.
-* Adversarial Robustness (resilience to adversarial attacks) [102], * System Reliability (overall system uptime and fault tolerance) [103]
-Table 2: Factors and Subfactors of Human-AI Collaboration with Associated Metrics for Evaluation
-Goals
-The Goals factor ensures that the collaboration has a clear direction and that both human and AI efforts are aligned towards shared outcomes. This factor encompasses two key subfactors:
-  • Individual Goals: The specific aims of each participant.
-For the AI, this could be learning efficiently or maximizing accuracy.
-For humans, it might involve providing effective teaching or achieving task objectives.
-Misalignment of individual goals can hinder collaboration, as highlighted in studies on migrating to HAIC, which emphasize the importance of goal congruence [40].
-    • Learning Curve: Measures the rate of improvement over time, indicating the effectiveness of human teaching and AI learning.
-This is plotted by graphing the performance of the AI (like accuracy) against the number of training samples or training iterations.
-A steep learning curve suggests the AI is learning effectively [81].
-    • Prediction Accuracy: The proportion of correct predictions made by the AI model out of all predictions.
-This can be measured using metrics like accuracy, F1 score, or area under the ROC curve, depending on the nature of the task [86].
-    • Teaching Efficiency: The gain in AI performance per unit of teaching effort by the human.
-This could be measured by tracking improvements in the AI’s accuracy or error rate after receiving human input.
-This concept is explored in works like that of Cakmak and Thomaz [85]
-  • Collective Goals: Shared objectives that maximize the overall performance of the HAIC system.
-For example, in a medical diagnosis system, the collective goal might be to improve patient outcomes through accurate and timely diagnosis.
-These collective goals are often broader than individual goals and necessitate the combined efforts of both humans and AI.
-    • Overall System Accuracy: It can be calculated as a weighted average of different performance metrics, reflecting the relative importance of each task in achieving the overall goal.
-Evaluates the general accuracy of the AI-human collaborative system, tailored to the specific objectives of the system.
-[65].
-    • Objective Fulfillment Rate: The percentage of predefined goals that are successfully achieved within the HAIC collaboration.
-This metric is often used in project management contexts to track progress towards overall objectives and can be adapted to assess the success of HAIC collaborations [87].
+* Ease of Use (user satisfaction with communication methods), * Clarity of Communication (understandability of queries and responses) [82, 83]
+Feedback Mechanisms Human feedback on AI queries, AI adjustments based on feedback.
+* Frequency of Feedback (number of feedback instances), * Feedback Quality (relevance and specificity of feedback), * Feedback Impact (measured change in AI performance due to feedback) [83, 84, 82]
+Adaptability AI’s ability to adapt query strategy and learning based on human input.
+* Strategy Adaptation Effectiveness (improvement in query selection based on feedback), * Adaptability Score (composite metric of flexibility and responsiveness) [82, 85, 86]
+Trust and Safety Trust in AI system, ensuring user safety.
+* Trust Score (level of user trust in AI system) [87], * Safety Incidents (number of safety-related issues) [88], * Error Reduction Rate (decrease in errors due to AI assistance) [2], * Confidence (level of user trust in AI recommendations) [89]
+Task Allocation Complementarity AI handles data analysis and uncertainty estimation; Humans provide expertise and clarification.
+* Query Efficiency (amount of information gained per query), * Expertise Utilization (degree to which human expertise is leveraged) [78]
+Flexibility Adjusting query strategy based on interaction dynamics.
+* Adaptability Score (same as in Interaction), * Dynamic Task Allocation (ability to adjust task distribution based on real-time needs) [23]
+Efficiency Streamlining the learning process through targeted queries.
+* Model Improvement Rate (rate of improvement in AI performance) [59, 78], * Learning Curve (same as in Goals), * Resource Utilization (efficiency in using computational resources) [59]
+Responsiveness AI’s ability to rapidly adjust to new inputs and feedback.
+* Response Time (time taken to respond to queries or feedback), * Real-time Performance (ability to handle time-sensitive tasks) [82, 90]
+Collaborative Decision Making Joint decisions on data labeling and correction of AI’s model.
+* Impact of Corrections (measured improvement in system performance through human intervention) [91, 92], * Decision Effectiveness (impact of decisions on overall task success) [93, 94]
+Continuous Learning AI’s ability to learn from human input and adjust its model.
+* Model Improvement Rate (same as in Efficiency), * Learning Curve (same as in Goals), * Knowledge Retention (ability to retain learned information over time) [95]
+Mutual Support AI’s ability to provide support to humans in the learning process; Humans support by providing information.
+* Error Reduction Rate (decrease in errors due to AI assistance) [2], * Confidence (level of user trust in AI recommendations) [96], * Task Completion Time (with and without AI support) [97]
+Robustness Ensuring AI system’s robustness and reliability. * Adversarial Robustness (resilience to adversarial attacks) [98], * System Reliability (overall system uptime and fault tolerance) [99]
+Goals Factor
+The Goals factor ensures that the collaboration has a clear direction and that both human and AI efforts are aligned towards shared outcomes.
+This factor encompasses two key subfactors: Individual Goals, which pertain to the specific aims of each participant, and Collective Goals, which represent shared objectives that maximize the overall performance of the HAIC system.
+Table 3 below summarizes the metrics used to evaluate these subfactors, providing concise definitions, measurement methods, and relevant references to support the evaluation of the Goals factor in Human-AI Collaboration (HAIC) systems.
+Table 3: Summary of metrics for the Goals factor in Human-AI Collaboration.
+Subfactor Metric Definition and Measurement
+Individual Goals Learning Curve Measures the rate of improvement over time, indicating the effectiveness of human teaching and AI learning.
+This is plotted by graphing AI performance (e.g., accuracy) over iterations or training samples.
+A steep learning curve suggests effective learning [76].
+Prediction Accuracy The proportion of correct predictions made by the AI model out of all predictions.
+This can be evaluated using metrics like accuracy, F1 score, or area under the ROC curve, depending on the task [80].
+Teaching Efficiency The gain in AI performance per unit of teaching effort by the human.
+This is measured by tracking improvements in AI accuracy or error rate after human input, as explored in studies such as Cakmak and Thomaz [79].
+Collective Goals Overall System Accuracy Calculated as a weighted average of performance metrics, reflecting the relative importance of each task in achieving the overall goal.
+This metric evaluates the general accuracy of the AI-human collaborative system, tailored to specific objectives [59].
+Objective Fulfillment Rate The percentage of predefined goals successfully achieved within the collaboration.
+Often used in project management contexts to track progress and adapted to assess HAIC success [81].
 Schneider et al.
@@
 [100] illustrate how human input influences AI’s learning processes, emphasizing the mutual dependence of individual and collective goals.
-Interaction
+Interaction Factor
 The Interaction factor underscores the critical communication mechanisms through which humans and AI systems exchange information, provide mutual feedback, and adaptively respond to each other’s inputs.
-Key subfactors within the Interaction factor include:
-  • Communication Methods: This involves evaluating the clarity, accessibility, and intuitiveness of the interfaces and protocols through which humans and AI interact.
-For instance, the clarity of communication in a medical diagnosis system might be assessed by how well the AI can explain its reasoning to the doctor in an understandable way [105, 90].
-Effective communication methods ensure efficient information exchange and mutual understanding, which can be evaluated using metrics such as:
-    • Clarity of Communication: Can be assessed through qualitative methods like expert review or user feedback, or through quantitative measures like the proportion of correctly interpreted queries or the time taken to resolve ambiguities.
-    • Ease of Use: Often measured through Likert scale questionnaires or user interviews to gauge the user’s subjective experience with the system.
-  • Feedback Mechanisms: These mechanisms assess how input from human operators is used to guide AI learning and behavior adjustments.
-For instance, in a customer service chatbot, feedback mechanisms could involve analyzing how user ratings or corrections influence the chatbot’s future responses.
-It’s crucial that AI systems can not only receive but also meaningfully integrate feedback to enhance their interactions with humans [83, 88].
-The effectiveness of feedback mechanisms can be measured through:
-    • Frequency of Feedback (number of feedback instances): This is a simple count of how often users provide feedback, indicating engagement and potentiallearning opportunities for the AI, as highlighted in Amershi et al.’s (2014) [83] work on power to the people.
-    • Feedback Quality (relevance and specificity of feedback): This can be measured qualitatively through expert review or quantitatively by analyzing the correlation between feedback and subsequent AI performance improvements as suggested by Dautenhahn (2007) [88] in the context of socially intelligent robots.
-    • Feedback Impact (measured change in AI performance due to feedback): This quantifies the tangible effect of feedback on AI behavior or output quality, often measured by improvements in accuracy, precision, or other relevant metrics.
-  • Adaptability: This measures the AI’s ability to modify its behaviors based on feedback and changes in the environment or task conditions.
-Adaptability is essential for AI systems to remain effective as collaboration needs evolve [106, 107].
-For example, in an autonomous vehicle, the AI system should be able to adapt its driving style based on the driver’s feedback or changes in traffic conditions.
-Adaptability can be evaluated through:
-    • Adaptability Score (composite metric of flexibility and responsiveness): This metric combines measures of how quickly and effectively the AI responds to changes and how flexibly it adjusts its strategies, which can be derived from studies like that of El-Assady et al.
-(2022) [90] on communication biases in HAIC.
-    • Strategy Adaptation Effectiveness (improvement in query selection based on feedback): In systems that involve querying for information, this metric measures whether the AI learns to ask more effective questions based on human responses, potentially incorporating insights from research on semantic interaction challenges by Wenskovitch et al.
-(2020) [89].
-  • Trust and Safety: Ensuring trust in the AI system and the safety of users is critical for effective HAIC. Metrics to evaluate trust and safety include:
-    • Trust Score: The level of user trust in the AI system, often measured through surveys and trust scales [100].
-    • Safety Incidents: The number of safety-related issues encountered, used to assess the reliability and safety of the AI system.
+Key subfactors within this factor include Communication Methods, Feedback Mechanisms, Adaptability, and Trust and Safety.
+These subfactors capture the essential elements of interaction in Human-AI Collaboration (HAIC), ensuring seamless and effective communication, the incorporation of feedback, and adaptability to dynamic conditions.
+Table 4 summarizes the metrics used to evaluate the Interaction factor, providing concise definitions and measurement methods for each subfactor.
+Table 4: Summary of metrics for the Interaction factor in Human-AI Collaboration.
+Subfactor Metric Definition and Measurement
+Communication Methods Clarity of Communication Assesses how understandable the communication is between humans and AI.
+This can be evaluated using qualitative methods like expert review or user feedback and quantitative measures like the proportion of correctly interpreted queries or time taken to resolve ambiguities [101, 86].
+Ease of Use Often measured through Likert scale questionnaires or user interviews to gauge the user’s subjective experience with the system.
+Feedback Mechanisms Frequency of Feedback Monitors how often users provide feedback, helping developers assess engagement levels and identify areas for improvement [102].
+Feedback Quality Measures the relevance and specificity of feedback, either qualitatively through expert reviews or quantitatively by analyzing the correlation between feedback and subsequent AI performance improvements.
+Hein et al.
+[103] explore feedback acceptance and motivation when provided by AI versus human supervisors.
+Feedback Impact Quantifies the tangible effect of feedback on AI behavior or output quality, often measured by improvements in accuracy, precision, or other relevant metrics [104].
+Adaptability Adaptability Score Combines measures of flexibility and responsiveness to assess how quickly and effectively the AI adapts to changes, as studied by El-Assady et al. [86].
+Strategy Adaptation Effectiveness Measures improvement in AI’s query selection based on human feedback, incorporating insights from semantic interaction research [85].
+Trust and Safety Trust Score Assesses user trust in the AI system through surveys and trust scales [105].
+Safety Incidents Tracks the number of safety-related issues encountered to evaluate the reliability and safety of the AI system.
 Ethical considerations such as bias, fairness, and transparency are critical to the success of HAIC.
@@
 [107] emphasize the importance of fairness in AI design and implementation, advocating for continuous monitoring and evaluation to mitigate bias.
-Task Allocation
+Task Allocation Factor
 Task allocation in HAIC focuses on strategically distributing responsibilities between humans and AI to leverage the strengths of each party and optimize overall performance.
 Effective task allocation enhances efficiency, minimizes errors, and increases user satisfaction, making it a critical component in the success of HAIC systems [108].
-Key subfactors in Task Allocation, and its relevant metrics include:
-  • Complementarity: Aligning tasks with the unique capabilities of humans and AI. For instance, AI’s analytical strengths can complement human decision making and creativity [111].
-    • Query Efficiency: Indicates the number of queries the AI needs to reach a certain level of accuracy, with fewer queries signifying more efficient learning [84].
-    • Expertise Utilization: The degree to which human expertise is leveraged, often measured qualitatively through observation or interviews, or quantitatively through metrics like the percentage of decisions where human input is decisive [26].
-  • Flexibility: The ability to adjust task allocation dynamically in response to changing circumstances or new challenges is crucial for maintaining operational effectiveness [112].
-Research by Dubey et al.
-(2020) [3] explores how AI can be used to optimize task allocation for improved efficiency.
-    • Dynamic Task Allocation: The ability to adjust task distribution based on real-time needs, assessed by how quickly and effectively the system adapts to changes in workload, priorities, or available resources, as explored in studies on human-robot collaboration by Tsarouchi et al.
-(2017) [112].
-  • Efficiency: Focuses on maximizing productivity and resource utilization by optimizing task distribution, often allowing AI to automate routine tasks [111, 39]. Efficiency can be measured by:
-    • Model Improvement Rate: The rate of improvement in AI performance, calculated by taking the difference between the AI’s performance at a later time point and its performance at an earlier time point.
-This difference is then divided by the time interval between these two points.
-This can be applied to various AI performance metrics like accuracy or error rates, and its significance in streamlining HAIC has been explored by studies such as Vossing et al.
-(2022) [65].
-    • Resource Utilization: The efficiency in using computational resources like CPU time, memory, or network bandwidth.
-This can be measured as a percentage of utilized resources compared to the total available, and can be a key indicator of overall system efficiency.
-    • Learning Curve: The rate of improvement in AI performance over time, as defined earlier in the Goals section.
-  • Responsiveness: The capability of systems to rapidly adjust to new requirements or unexpected situations is essential for maintaining workflow continuity, as discussed by Johannsmeier and Haddadin (2016) [113].
-Responsiveness can be evaluated through:
-    • Response Time: The average time taken for the AI or human to respond to queries or feedback.
-A lower response time indicates quicker and more effective communication, as highlighted in studies on human-robot interaction by Hemmer et al.
-(2022) [111, 82].
-    • Real-time Performance: The ability to handle time-sensitive tasks, measured by factors like throughput (tasks completed per unit time) and latency (delay in response) [94].
-This is particularly important in applications like real-time decision support systems and autonomous vehicles.
-    • Learning Curve: The rate of improvement in AI performance over time, as defined earlier in the Goals section.
-  • Collaborative Decision Making [111]: Involves integrating the insights and expertise of both humans and AI to make informed decisions, ensuring that all parties contribute to problem- solving.
-    • Impact of Corrections: Measured improvement in system performance resulting from human intervention, often quantified by changes in accuracy or error rates after corrections are made.
-This can be used to assess the effectiveness of human input in refining AI outputs [96, 95].
-    • Decision Effectiveness: The impact of decisions on the overall task success, measured by comparing outcomes of decisions made with and without AI collaboration.
-This metric can be customized based on the specific goals of the HAIC system [98, 97].
-  • Continuous Learning: AI systems should evolve by learning from past interactions to enhance future performance and adaptation.
-Yin et al.
-(2021) highlight the importance of continuous learning for optimizing task allocation in dynamic environments [114].
-    • Knowledge Retention: The AI’s ability to retain learned information over time, often assessed by testing the AI’s performance on tasks after a period of non-use.
-This metric is important for understanding the long-term impact of human feedback and training on the AI system.
-    • Model Improvement Rate: (same as in Efficiency)
-    • Learning Curve: (same as in Goals)
-  • Mutual Support: Promoting a supportive environment where both humans and AI can offer reciprocalassistance enhances the overall collaboration effectiveness [111].
-Mutual support can be evaluated through:
-    • Error Reduction Rate: Measures the decrease in errors or mistakes made by the AI following human intervention, showing the effectiveness of human corrections  [4].
-    • Confidence: The level of user trust in AI recommendations, often measured using Likert-scale questionnaires or surveys [100].
-Higher confidence scores indicate a stronger belief in the AI’s capabilities and a greater likelihood of accepting its recommendations.
-    • Task Completion Time: Comparison of task completion times with and without AI support [101]. This metric helps assess the efficiency gains resulting from HAIC.
-  • Robustness: Ensuring the AI system performs reliably under various conditions, including adversarial attacks and different domains, is crucial for HAIC effectiveness.
-Robustness can be evaluated through:
-    • Adversarial Robustness: The resilience of the AI system to adversarial attacks, measured through testing against adversarial examples.
-    • System Reliability: The overall system uptime and fault tolerance, measured through system logs and reliability tests.
+Table 5 summarizes the key subfactors and metrics associated with task allocation, providing concise definitions and measurement methods.
+Table 5: Summary of metrics for the Task Allocation factor in Human-AI Collaboration.
+Subfactor Metric Definition and Measurement
+Complementarity Query Efficiency Indicates the number of queries the AI needs to reach a certain level of accuracy. Fewer queries signify more efficient learning [78].
+Expertise Utilization The degree to which human expertise is leveraged.
+Measured qualitatively through observation or interviews, or quantitatively through metrics like the percentage of decisions where human input is decisive [19].
+Flexibility Dynamic Task Allocation The ability to adjust task distribution based on real-time needs.
+Assessed by how quickly and effectively the system adapts to changes in workload, priorities, or resources [109].
+Efficiency Model Improvement Rate The rate of improvement in AI performance.
+Calculated as the difference in performance between two time points, divided by the time interval.
+Applied to metrics like accuracy or error rates [59].
+Resource Utilization The efficiency of computational resource use (e.g., CPU time, memory). Measured as a percentage of utilized resources compared to the total available.
+Learning Curve The rate of improvement in AI performance over time, as previously defined in the Goals section.
+Responsiveness Response Time The average time taken for AI or human participants to respond to queries or feedback. A lower response time indicates quicker and more effective communication [110, 82].
+Real-time Performance The ability to handle time-sensitive tasks, measured by factors like throughput (tasks completed per unit time) and latency (delay in response) [90].
+Learning Curve The rate of improvement in AI performance over time, as defined earlier.
+Collaborative Decision Making Impact of Corrections Measures the improvement in system performance from human intervention. Quantified by changes in accuracy or error rates after corrections [92, 91].
+Decision Effectiveness The impact of decisions on task success. Measured by comparing outcomes of decisions made with and without AI collaboration [94, 93].
+Continuous Learning Knowledge Retention Assesses the AI’s ability to retain learned information over time. Evaluated by testing performance after periods of non-use [111].
+Model Improvement Rate Same as defined in the Efficiency section.
+Learning Curve Same as defined in the Goals section.
+Mutual Support Error Reduction Rate Measures the decrease in errors made by the AI following human intervention. Demonstrates the effectiveness of human corrections [2].
+Confidence The level of user trust in AI recommendations. Measured using Likert-scale questionnaires or surveys [105].
+Task Completion Time Compares task completion times with and without AI support. Helps assess efficiency gains from HAIC [97].
+Robustness Adversarial Robustness The resilience of the AI system to adversarial attacks. Measured through testing against adversarial examples.
+System Reliability The overall system uptime and fault tolerance. Evaluated through system logs and reliability tests.
 This structured approach, detailing subfactors and metrics, ensures a comprehensive evaluation of HAIC systems, enabling continuous improvement and effective collaboration between humans and AI.
@@
 These metrics are essential for quantitatively evaluating the efficacy, efficiency, and adaptability of HAIC systems across multiple domains.
-{adjustwidth}
+Table 6: Evaluation Metrics for Human-AI Collaboration
+Metric Formula/Description Example
+Learning Curve Graphical representation of performance vs.
+number of training samples.
+In a medical image analysis task, a graph shows the AI’s accuracy in diagnosing tumors steadily increasing over 1000 training images.
 […]
-Table 3: Evaluation Metrics for Human-AI Collaboration
+[…]
+[…]
+[…]
+Feedback Impact Impact = $\text{Performance Post-Feedback}-\text{Performance Pre-Feedback}Performance Post-Feedback - Performance Pre-Feedback$ A language translation AI’s BLEU score improves from 0.65 to 0.72 after incorporating user feedback on mistranslations.
+Adaptability Score Adaptability Score = $\text{Performance Post-Adaptation}-\text{Performance Pre-Adaptation}Performance Post-Adaptation - Performance Pre-Adaptation$ A self-driving car adapts its driving style based on weather conditions, improving safety scores by 12%.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Impact of Corrections Improvement = $\text{Performance Post-Correction}-\text{Performance Pre-Correction}Performance Post-Correction - Performance Pre-Correction$ Human corrections to an AI writing assistant’s suggestions result in a 15% improvement in the text’s readability score.
+[…]
+[…]
+Task Completion Time Time Saved = Time Without $AI\text{Time Without AI}-\text{Time With AI}Time With$out AI - Time With AI An AI-powered manufacturing robot reduces the time to complete a specific task from 10 minutes to 5 minutes.
+[…]
+Safety Incidents Safety Incidents = Number of safety-related issues encountered A self-driving car system logs 2 safety incidents per 1000 miles driven.
+[…]
+[…]
+[…]
 3.2 Building the Framework & Quantifying the Evaluation of HAIC
@@
 By providing a structured and comprehensive evaluation approach, this framework aims to enhance the understanding and assessment of HAIC, fostering the development of more effective and collaborative human-AI systems.
+Each metric is color-coded based on its primary factor (Table 2), enhancing clarity and usability.
+This adaptable, scalable framework fosters a comprehensive understanding of HAIC systems, improving their assessment and development.
 AI-Centric Collaboration
-Questions in this mode emphasize the AI’s autonomous capabilities and performance.
-We focus on efficiency metrics (like Model Improvement Rate) and uncertainty handling, as these are core challenges for AI systems operating primarily independently.
-The AI-Centric mode is particularly relevant for systems where AI plays a dominant role in decision-making, such as predictive analytics and automated systems.
-How efficiently does the AI make decisions independently?
-  • Learning Curve: Measures the rate of improvement in AI performance over time.
-  • Model Improvement Rate: Measures improvement over time, indicating decision-making efficiency.
-  • Query Efficiency: Evaluates how much information is gained with minimal queries.
-  • Resource Utilization: Assesses how well the AI manages computing power.
-  • Response Time: Tracks how quickly the AI makes decisions.
-  • Real-time Performance: Evaluates the AI’s ability to handle time-sensitive tasks.
-  • Objective Fulfillment Rate: Measures the percentage of tasks successfully completed.
-How well does the AI handle uncertainty?
-  • System Accuracy: Measures overall performance, partially reflecting uncertainty handling.
-  • Error Reduction Rate: Indicates if the AI makes fewer mistakes over time.
-  • Prediction Accuracy: Quantifies correct outputs specifically in uncertain scenarios.
-  • Feedback Impact: Shows how much human guidance improves AI decisions under uncertainty.
+Focuses on AI autonomy, emphasizing efficiency and uncertainty handling metrics.
 Human-Centric Collaboration
-Here, the questions center on how well the AI communicates with and supports the human user.
-Metrics like Clarity of Communication and Ease of Use are essential for a positive user experience.
-The Human-Centric mode is particularly relevant for systems where human input is crucial for decision-making, such as in creative fields or complex problem-solving.
-How clear and usable is the AI’s communication?
-  • Clarity of Communication: Directly assesses if queries and responses are understandable.
-  • Ease of Use: Measures overall user satisfaction with the interaction.
-  • User Satisfaction: Indicates if the AI’s communication methods are acceptable to humans.
-  • Frequency of Feedback: Quantifies the number of feedback instances.
-How effectively does the AI support human tasks?
-  • Teaching Efficiency: Gauges how much effort is needed to train the AI.
-  • Error Reduction Rate: Quantifies how much the AI helps in reducing human errors.
-  • Confidence: Measures user trust in the AI’s recommendations.
-  • Task Completion Time: Indicates if AI assistance makes tasks faster.
+Evaluates AI’s communication and support for human tasks, emphasizing usability and user satisfaction.
 Symbiotic Collaboration
-This mode’s questions strike a balance between adaptability and task allocation.
-We ask about the system’s flexibility in response to feedback (Adaptability Score) and the effectiveness of resource distribution, reflecting the collaborative core of Symbiosis.
-How easily does the AI adapt to human feedback?
-  • Adaptability Score: A composite metric reflecting flexibility and responsiveness.
-  • Strategy Adaptation Effectiveness: Measures how well the AI changes approaches based on input.
-  • Feedback Quality: Assesses how valuable the provided human feedback is.
-  • Feedback Impact: Shows the improvement resulting from human guidance.
-How well are tasks allocated to optimize both human and AI strengths?
-  • Dynamic Task Allocation: Reflects the ability to shift task distribution in real-time.
-  • Query Efficiency: Assesses how well information is requested to leverage human expertise.
-  • Expertise Utilization: Measures the degree to which human skills are leveraged.
-  • Knowledge Retention: Evaluates if the AI learns effectively from human input.
-  • Model Improvement Rate: Measures how much human input contributes to AI progress.
-  • Decision Effectiveness: Indicates the impact of decisions on overall task success.
-Advantages of the Decision Tree Approach
-This methodology offers several benefits:
-  • User-Friendly: Simplifies the evaluation process, making it accessible to users with varying levels of expertise.
-  • Tailored Assessment: Ensures that the evaluation is relevant to the specific characteristics and goals of the HAIC system.
-  • Adaptable Framework: Accommodates the evolving nature of HAIC technologies and practices.
-Although the decision-tree based framework offers a solid foundation for evaluating HAIC, the field continues to make significant advancements.
-As AI systems advance and expand their range of uses, various challenges and prospects arise.
-In order to maintain effectiveness, it is necessary for our evaluation methods to also adjust and change.
+Assesses adaptability and task distribution to optimize human-AI strengths.
+By structuring evaluation through this decision-tree approach, we enhance consistency, enabling better-informed HAIC system development and optimization.
+This approach simplifies evaluation, tailors assessments to HAIC systems, and remains adaptable as technologies evolve.
 Weighting Mechanism & Overall Score Calculation
 While the decision tree and associated metrics lead the evaluation process, we recognize the importance of a weighting mechanism to provide a more comprehensive evaluation of HAIC effectiveness.
-This system would weight each metric according to its relevance to the specific HAIC mode and the evaluation’s criteria.
-For example, in an AI-Centric system, accuracy might be assigned a higher weight than user satisfaction.
-Conversely, in a Human-Centric system, ease of use and trust could be prioritized.
-In the Symbiotic mode, the weights might be distributed more evenly across metrics reflecting collaboration and mutual benefit.
-The overall score for a HAIC system could then be calculated as the weighted average of the individual metric scores.
-However, determining appropriate weights requires further research and validation across different real - world paradigms.
-Several studies in the literature have utilized weighting mechanisms to evaluate complex systems.
-For example, Kamar et al.
-(2012) [116] employed a weighted approach to combine human and machine contributions in a crowdsourcing system, emphasizing the importance of weighting in achieving balanced evaluations.
-Similarly, Amershi et al.
-(2014) [83] discussed the role of weighting different interaction metrics to assess the effectiveness of human-AI collaboration tools, highlighting how different priorities can shift the weighting schema.
-Moreover, Vossing et al.
-(2022) [65] implemented a weighting mechanism to balance transparency and performance in AI systems, demonstrating that context-specific weighting can enhance the evaluation’s relevance and accuracy.
-These examples underscore the practical value of weighting mechanisms and support their inclusion in our proposed framework.
-By incorporating a validated weighting mechanism, our framework ensures a nuanced and context-sensitive assessment of HAIC systems, aligning with best practices in the literature.
-Future research should focus on refining these weights through empirical studies across various HAIC applications to enhance the robustness and applicability of the evaluation framework.
+A weighting mechanism ensures a comprehensive evaluation, adjusting metric importance based on HAIC mode.
+Weighted averages determine overall scores, enhancing accuracy and adaptability.
+Prior research supports this approach, emphasizing context-specific evaluation [113][83][59].
+Future work should refine these weights through empirical studies to strengthen HAIC assessment robustness.
 4 Implications of the HAIC Evaluation Framework to Real-World Domains
+Although it has not yet been empirically validated, the proposed HAIC framework offers theoritical basis for assesing human-AI collaboration across sectors such as manufacturing, healthcare, finance, and education.
+By emphasizing factors like Adaptability, Confindence and Clarity of Communication this framework lays the groundwork for future empirical studies.
 4.1 Manufacturing
-The manufacturing industry’s focus on safety, accuracy, and productivity closely corresponds with the Symbiotic mode of HAIC, which requires the collaboration of human expertise and AI skills to get the best results.
-This mode prioritizes metrics like Adaptability Score and Confidence.
-Recent study highlights a shift from focusing primarily on productivity- centered measures to a more worker-centric approach, placing greater emphasis on safety and trust.
-Research such as Rabby et al.’s (2020) [117] time-driven performance-aware trust model and Okamura and Yamada’s (2020)  [118] adaptive trust calibration approach exemplify this change.
-These works emphasize the importance of the Interaction factor, specifically focusing on the subfactors of Communication Methods and Adaptability.
-They achieve this by creating systems that can adapt to human actions and uncertainties.
-The primary emphasis of Amin et al.’s (2020)  [119] mixed-perception method is on safety, which aligns with the Task Allocation factor’s concentration on Complementarity and Mutual Support.
-This strategy involves the collaboration between human and AI systems to promote worker protection.
-The focus placed by Ming Hou (2021)  [120] on context-based interface design provides additional support for the Symbiotic method.
-Our methodology proposes the following essential criteria for evaluating HAIC systems in manufacturing: Adaptability Score, Error Reduction Rate, Confidence, and Task Completion Time.
-These measurements will assess the systems’ capacity to adapt to human operators, enhance safety, foster worker confidence, and ultimately optimize industrial processes.
+Manufacturing’s focus on safety, accuracy, and productivity aligns with the HAIC Symbiotic mode, which integrates human expertise with AI skills.
+Studies [114, 16, 115, 116] demonstrate a shift towards worker-centric, adaptive approaches that highlight critical metrics such as Adaptability Score, Error Reduction Rate, Confidence, and Task Completion Time.
 4.2 Healthcare
-The integration of AI in healthcare, particularly in medical imaging, often operates in a collaborative mode, where AI augments human expertise to achieve more accurate diagnoses.
-Although AI is highly effective in medical image classification and segmentation  [121], human skill is still essential for the interpretation and ultimate diagnosis.
-Contributions to this collaborative approach are made by studies such as Zheng et al.’s (2023)) [122], Eldin & Kaboudan’s (2023) [123] study of AI-driven platforms, and Aleid et al.’s (2023) [53] work on tumor identification.
-These developments specifically focus on the Goals aspect of the HAIC framework, with a special emphasis on enhancing diagnostic Accuracy and Efficiency.
-The Communication Methods subfactor is crucial as AI systems must provide unambiguous and comprehensible outputs, such as the Clarity of Communication.
-In the ICU setting, as explored by Dziorny et al.
-(2022) [124], the quality of Decision Support provided by AI becomes crucial, emphasizing the need for evaluating metrics like Decision Effectiveness.
-Our approach emphasizes key parameters such as System Accuracy, Prediction Accuracy, Response Time, and Clarity of Communication to successfully evaluate these collaborative HAIC systems.
-Future study should investigate the impact of such systems on the workflow of radiologists and oncologists, specifically in relation to Decision Support and Time-to-Diagnosis.
-Furthermore, conducting research to examine any possible effects on Patient Communication would be important.
+In medical imaging diagnosis, AI supports human decision-making by enhancing diagnostic accuracy and efficiency [117, 118, 119, 47, 120].
+The framework targets improvements in diagnostic Accuracy and Efficiency, emphasizing clear AI outputs.
+Evaluation parameters include System Accuracy, Prediction Accuracy, Response Time, and Decision Effectiveness, with further research needed on workflow impacts and patient communication.
 4.3 Finance
-The finance sector largely employs HAIC in a Symbiotic manner, leveraging AI’s analytical capabilities alongside human financial experience.
-This collaboration is evident in the work of Atadoga et al.
-(2024) [125], Al-Fatlawi et al.
-(2024)  [54], Jain (2023) [126], and Ali et al.
-(2022) [127].
-These studies highlight the importance of Goals such as Accuracy and Efficiency in minimizing fraud and operational costs.
-However, the Symbiotic nature of this collaboration also emphasizes the Interaction factor, specifically the subfactor of Clarity of Communication, as AI-generated insights must be clearly communicated to human decision-makers.The Task Allocation factor is evident in the division of labor, with AI systems often identifying potential fraud while human experts analyze complex cases.
-The Error Reduction Rate is a key metric to assess the effectiveness of AI in this context.
-Future research should investigate the ethical and security implications of these systems, along with their long-term impacts on client relationships and the quality of financial decisions.
+The finance sector benefits from a Symbiotic approach where AI’s analytical capabilities complement human expertise to enhance fraud detection and operational efficiency [121, 48, 122, 123].
+Key aspects include achieving goals of Accuracy and Efficiency and ensuring clear communication of AI insights.
+Metrics such as the Error Reduction Rate help assess system effectiveness, with future work needed to address ethical and long-term impacts.
 4.4 Education
-HAIC in Education demonstrates a hybrid of Symbiotic and Human-Centric paradigms, with AI playing a supportive role in enhancing both teaching practices and learner experiences.
-Studies by Ifenthaler and Schumacher (2023) [128], Ji et al.
-(2023) [129], Wang et al.
-(2020)  [40], and Sharma et al.
-(2023) [49] highlight these modes.
-These works target Goals such as enhancing Teaching Efficiency and achieving personalized learning through Individualized Goals.
-The Interaction factor is crucial, as effective communication between AI tutors and learners impacts the learning process, necessitating assessment of Clarity of Communication and Ease of Use.
-The Lumilo project, investigated by Holstein and Aleven (2022) [12], shows Symbiotic collaboration for improved learning outcomes, where real-time analytics provided by AI enhance the teacher’s ability to provide Mutual Support to students.
-Evaluating the impact of these systems requires metrics like Task Completion Time and Learning Curve to ensure a comprehensive understanding of both student and teacher experiences.
-Further research should emphasize fostering student agency and critical thinking within HAIC frameworks, moving beyond passive consumption of AI-generated tools toward active engagement and co-creation of knowledge, as advocated by Fadiya (2023)  [130].
-6 The Unique Challenge of Evaluating Creative and Linguistic AI
-As seen in the previous sections, establishing comprehensive frameworks for assessing HAIC is vital, but certain applications pose distinct challenges due to their unique nature.
-Two prominent examples are Large Language Models (LLMs) and Generative AI in the Arts, where AI’s impact on human creativity, expression, and communication requires specialized evaluation methodologies.
-While our existing framework provides a foundation for HAIC assessment, these creative and linguistic domains raise questions about the adequacy of traditional metrics.
-Can quantitative measures, such as the ones listed in Table 3 fully capture the nuances of artistic collaboration or the impact of AI on human language?
-How can we evaluate the ethical implications of AI- generated art or the potential for bias in LLMs?
-These challenges necessitate a more nuanced and tailored approach to evaluation, as we will explore in the following sections.
-6.1 Large Language Models (LLMs)
-Large Language Models (LLMs) have revolutionized text generation, translation, and dialogue.
-The GPT-3 [131] family, including ChatGPT and GPT-4, showcases the power of these models to engage in complex natural language processing [132].
-While their potential for transforming Human-AI Collaboration (HAIC) is vast, evaluating these interactions requires a critical lens and the development of new methodologies [133].
-Ensuring the interpretability and fairness of LLMs is crucial.
-Binz & Schulz’s (2022) [134] study reveals that even when LLMs perform well on cognitive tasks, they might employ reasoning patterns that are opaque or reflect biases embedded in their training data.
-Evaluation frameworks must incorporate methods for assessing how well LLMs explain their outputs and whether biases are mitigated over time.
-Additionally, the quality of human interaction with LLMs is paramount.
-Using GPT-3 to simulate human interactions, as in the work by Aher, Arriaga, & Kalai (2022) [56], demonstrates the value of testing how well LLMs adapt to different users and communication styles.
-LLMs promise to augment cognitive tasks, enhance creativity, and personalize learning.
-Webb, Holyoak, & Lu (2022) [135] demonstrate an LLM’s capability for analogical reasoning, a core element of problem-solving.
-Further research should evaluate how such abilities impact HAIC success in real-world tasks.
-Regarding creativity, Douglas (2023) [136] highlights LLMs’ transformative role, inviting us to develop metrics for assessing how these tools amplify human ideation and artistic expression.
-In education, Liu et al.
-(2023) [137] explore the potential of ChatGPT and GPT-4 models to tailor learning experiences.
-Evaluating their effectiveness in this context will necessitate student-centered metrics focused on knowledge gains and learner autonomy.
-Longitudinal studies are vital for a nuanced understanding of LLM-based HAIC.
-Work by Yin et al.
-(2023) [138] on LLM self-knowledge suggests ways to track how both the AI’s capabilities and the user’s trust in the system evolve over extended periods of interaction.
-The work of Dwivedi et al.
-(2023) [139] highlights crucial challenges posed by ChatGPT, reminding us that even with the potential benefits, addressing societal implications and ethical use is paramount.
-The development of robust evaluation frameworks for LLM-powered HAIC is a rapidly evolving field.
-By addressing interpretability, bias, the quality of interaction, and the transformative potential of LLMs on human cognition, such frameworks will drive the creation of responsible, effective, and truly collaborative AI systems.
-6.2 Generative AI in the Arts
-Generative AI is revolutionizing artistic expression by enabling algorithms to generate artworks in various modalities, including visual art, music, and poetry.
-This technology transforms AI systems from tools into active co-creators, expanding the boundaries of what artists can achieve.
-Notable examples include AI-generated paintings, musical compositions, and even interactive installations tailored to audience responses.
-Studies like that of Gmeiner, Holstein, & Martelaro (2022) [140] suggest that understanding how effective human teams learn can provide a framework for designing generative AI systems that maximize collaborative potential.
-Evaluating HAIC in the context of generative AI for artists poses several challenges.
-Traditional art metrics such as aesthetic quality, originality, technical skill, audience reception, market value, and art historical significance provide a starting point for this evaluation [141].
-These metrics, which typically focus on the final artwork, can be adapted to also consider the collaborative process between the artist and the AI.
-Verheijden & Funk’s (2023) [57] work highlights the importance of assessing the generative AI’s impact on the design process itself - how it boosts inspiration, facilitates communication, and shapes the creative journey.
-The field necessitates methodologies that evaluate AI-assisted artistic creation holistically.
-Frameworks inspired by HAIC models developed for other domains will require adaptation to consider factors unique to artistic collaboration.
-He, Shrestha, Puranam, & Miron-Spektor (2023) [142] propose a novel theory of co-creativity as a joint search process, offering an insightful lens for analyzing the dynamic between human artist and generative AI system.
-The rise of AI-assisted art brings both ethical considerations and profound societal implications.
-Issues of copyright, ownership, and the meaning of authenticity in this new landscape demand attention.
-Epstein et al.
-(2023) [143] call for interdisciplinary investigation into generative AI’s cultural, economic, and legal impacts.
-Additionally, as AI tools become ubiquitous, reevaluating the definition of "artist" and determining the skills that foster success in this evolving domain become essential considerations.
-As generative AI in art continues to mature, its transformative potential is immense.
-We might witness a shift towards real-time, personalized art experiences, requiring HAIC evaluation models to be responsive and adaptive.
-Suh, Youngblom, Terry, & Cai (2021) [144] offer intriguing insights into how AI might reshape social interactions within creative settings.
-Moreover, focusing on documenting the artist-AI creative process itself could become as paramount as evaluating any final product.
+In education, a mix of Symbiotic and Human-Centric models enhances teaching and learning [124, 125, 34, 43, 10, 126].
+The focus is on improving Teaching Efficiency and achieving personalized learning through effective communication.
+Important metrics include Task Completion Time and Learning Curve, while further research should explore fostering active student engagement and critical thinking.
 5 Limitations in Investigating Behavioral Factors
-Our study provides a thorough framework to evaluate Human-AI Collaboration (HAIC) in different sectors, emphasizing goals, interaction dynamics, and task allocation.
-However, it does not explicitly examine behavioral aspects as a distinct main category of analysis.
-The exclusion of behavioral aspects of HAIC is not because they are unimportant but is due to methodological and scope concerns.
-The study of behavioral aspects in the context of HAIC involves a multidisciplinary approach, incorporating knowledge from psychology, sociology, cognitive science, and information technology to address its complexity.
-The complexity of the research requires interdisciplinary collaboration and skills outside the current study’s context.
-Behavioral aspects are important but can vary significantly and depend greatly on the context.
-The wide range of human behavior, shaped by cultural, psychological, and situational elements, presents a difficulty in establishing a uniform or universal evaluation framework for this topic.
-Extensive empirical study, including longitudinal studies and ethnographic methodologies, like those in Yang, Carroll, & Dragan (2022) [55], would provide a richer picture of how AI adaptability impacts human trust over time and how trust, in turn, shapes collaborative goals.
-Our focus on the structural elements of HAIC, like goals, interaction, and task allocation, enables a more objective examination based on observable and quantifiable results.
-Behavioral elements demand methodological techniques and analytical frameworks different from those used in our study due to their subjective interpretations and qualitative judgments.
-Behavioral variables are recognized as crucial for the success and efficacy of HAIC systems.
-We support more study focused on understanding these factors thoroughly.
-These studies could provide useful insights into the mutually beneficial connection between human views, beliefs, and behaviors and AI collaboration.
-This would help in creating more sophisticated, efficient, and human-centered HAIC systems.
-Behavioral aspects are crucial in understanding and optimizing HAIC.
-However, they were intentionally excluded as a primary emphasis in this study to keep the scope focused and manageable.
-We urge further study to fill this gap, expanding on the groundwork established by our approach to investigate the complex interaction between human behavior and AI collaboration.
+Our framework for evaluating Human-AI Collaboration (HAIC) emphasizes goals, interaction dynamics, and task allocation while excluding behavioral aspects due to methodological and scope constraints.
+Investigating behavioral factors involves interdisciplinary approach from psychology, sociology, cognitive science, which require complex and varried methodologies.
+While these factors are crucial for understanding trust and collaborative goals over time—exemplified by studies such as Yang, Carroll, & Dragan (2022) [49]—our study focuses on structural elements that are more readily quantifiable.
+We advocate for future research to explore behavioral dimensions in-depth to complement our objective framework.
 6 Conclusion
-This paper has explored the current state in the evaluation of HAIC and has illuminated a notable gap in the current HAIC evaluation methodologies: the absence of a framework that effectively integrates qualitative metrics alongside quantitative ones, applicable across a wide range of sectors.
-Our investigation revealed that existing evaluation practices are predominantly domain-specific, lacking the structured adaptability required for comprehensive HAIC assessment.
-The framework proposed herein endeavors to address this void by offering a structured approach to HAIC evaluation, one that is predicated on the mode of collaboration rather than the sector, application, or AI method used.
-This decomposition of evaluation into smaller, manageable components presents a novel way to scrutinize HAIC in its multifaceted nature.
-The significance of our framework lies in its universal applicability, affording a comparative lens to examine different collaboration methods.
-By transcending domain-specific limitations, it paves the way for a standardized evaluation approach that could be adopted across various sectors, enhancing the consistency and reliability of HAIC assessments.
-While the framework stands as a theoretical construct, its real-world implications remain untested.
-The future trajectory of our research is clear: to implement this framework as a software platform and apply it across diverse sectors and collaboration modes to validate its efficacy and adaptability.
-Such empirical endeavors will be crucial in refining the framework and ensuring its robustness.
-Our study acknowledges its limitations, particularly the omission of behavioral factors and ethical considerations from the framework.
-These elements, while outside our current scope, are critical dimensions of HAIC that warrant attention in future research.
-As we look to the horizon, our belief is steadfast in the potential of this framework to significantly advance the methods of evaluating HAIC.
-As this interdisciplinary field evolves, a standardized framework becomes increasingly crucial for its adoption from industry to society.
-We anticipate that our contribution will serve as a catalyst for future studies, spurring the development of HAIC evaluations that are as dynamic and intelligent as the collaborations they aim to measure.
+This paper makes a two-fold contribution to HAIC evaluation.
+First, it offers a comprehensive literature review of the existing methodologies and their limitations.
+Second, it introduces a structured framework that integrates key evaluation factors identified in the literature and provides a standardized method for assessing HAIC across various domains.
+A key gap in current methodologies is the absence of a framework that integrates both qualitative and quantitative metrics across diverse sectors.
+The framework proposed herein addresses this limitation by offering a structured evaluation approach based on the mode of collaboration rather than a specific sector or AI application.
+By breaking down evaluation into smaller, manageable components, it provides a flexible method to assess HAIC across different contexts.
+Although the framework is theoretical at this stage, future research will focus on empirical validation through software implementation and real-world applications across various sectors, ensuring its adaptability and robustness.
+Additionally, our study acknowledges limitations, particularly the exclusion of behavioral and ethical considerations from the evaluation framework.
+These dimensions, though outside the current scope, are essential aspects of HAIC that should be explored in future research.
 Acknowledgments
-The work leading to these results has received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No.
-101120218, project HumAIne (Hybrid Human-AI Decision Support for Enhanced Human Empowerment in Dynamic Situations).
+The work leading to these results has received funding from the EU’s Horizon Europe research and innovation programme under GA No.
+101120218, project HumAIne.
 Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union.
 Neither the European Union nor the granting authority can be held responsible for them.
+Appendix A The Unique Challenge of Evaluating Creative and Linguistic AI
+Establishing comprehensive frameworks for assessing HAIC is vital, however certain applications pose distinct challenges due to their unique nature.
+Two prominent examples are Large Language Models (LLMs) and Generative AI in the Arts, where AI’s impact on human creativity, expression, and communication requires specialized evaluation methodologies.
+While our framework provides a foundation, traditional metrics may be inadequate for these domains.
+Can quantitative measures fully capture artistic collaboration or AI’s influence on human language?
+How do we assess ethical concerns, bias, or AI-generated content’s societal impact?
+These challenges call for a nuanced evaluation approach.
+Although we don’t address these challenges in this work, we briefly discuss them here for completeness.
+A.1 Large Language Models (LLMs)
+LLMs have transformed text generation, translation, and dialogue.
+The GPT-3 [127] family, including ChatGPT and GPT-4, demonstrates their capacity for complex NLP tasks [128].
+Evaluating HAIC with LLMs requires developing methodologies beyond traditional metrics [129].
+Ensuring interpretability and fairness is vital.
+Binz & Schulz (2022)[130] show that even high-performing LLMs may use opaque reasoning patterns.
+Evaluation frameworks must assess how well LLMs explain their outputs and mitigate biases.
+Aher et al.
+(2022)[50] highlight the importance of testing LLM adaptability in user interactions.
+LLMs augment cognition, enhance creativity, and personalize learning.
+Webb et al.
+(2022)[131] show their potential in analogical reasoning.
+Douglas (2023)[132] emphasizes their impact on ideation and artistic expression, while Liu et al.
+(2023) [133] explore their role in education.
+Evaluating these aspects requires metrics for creativity, knowledge retention, and learner autonomy.
+Longitudinal studies provide insights into HAIC with LLMs.
+Yin et al.
+(2023)[134] suggest tracking how AI capabilities and user trust evolve over time.
+Dwivedi et al.
+(2023) [135] address societal implications, underscoring the need for ethical AI use.
+Developing robust LLM evaluation frameworks is essential. Addressing interpretability, bias, and interaction quality will help create responsible, effective HAIC systems.
+A.2 Generative AI in the Arts
+Generative AI is reshaping artistic expression across visual art, music, and poetry.
+AI moves beyond tools to co-creators, expanding artistic possibilities.
+Notable examples include AI-generated paintings, compositions, and interactive installations.
+Gmeiner et al.
+(2022) [136] suggest leveraging team learning frameworks to enhance AI collaboration in creative fields.
+Evaluating HAIC in generative AI requires adapting traditional art metrics—such as aesthetic quality, originality, and audience reception—to consider AI’s role in the creative process [137].
+Verheijden & Funk (2023) [51] stress assessing AI’s impact on ideation, communication, and artistic workflows.
+New methodologies must evaluate AI-assisted artistic creation holistically.
+He et al.
+(2023)[138] propose co-creativity as a joint search process, offering insights into human-AI dynamics.
+Ethical concerns also arise, including copyright, ownership, and authenticity.
+Epstein et al.
+(2023)[139] call for interdisciplinary research into AI’s cultural and economic impact.
+As AI-assisted art evolves, evaluation models must adapt.
+Future research should examine real-time, personalized experiences and how AI reshapes social interactions in creative settings [140].
+Documenting the artist-AI collaboration process may become as important as assessing final works.